### v1

# Explicit $7$-torsion in the Tate–Shafarevich groups of genus $2$ Jacobians

Abstract.
We describe an algorithm which, on input a genus $2$ curve $C/\mathbb{Q}$ whose Jacobian $J/\mathbb{Q}$ has real multiplication by a quadratic order in which $7$ splits, outputs twists of the Klein quartic curve parametrising elliptic curves whose mod $7$ Galois representations are isomorphic to a sub-representation of the mod $7$ Galois representation attached to $J/\mathbb{Q}$. Applying this algorithm to genus $2$ curves of small conductor in families of Bending and Elkies–Kumar we exhibit a number of genus $2$ Jacobians whose Tate–Shafarevich groups (unconditionally) contain a non-trivial element of order $7$ which is visible in an abelian three-fold.
1. Introduction
Let $K$ be a number field and let $A/K$ be an abelian variety. For each place $v$ of $K$ we denote the completion of $K$ at $v$ by $K_{v}$. We write $G_{K}=\operatorname{Gal}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu/K)$ for the absolute Galois group of $K$ and write $G_{v}=\operatorname{Gal}(\mkern 1.5mu\overline{\mkern-1.5muK_{v}\mkern-1.5mu}\mkern 1.5mu/K_{v})$. The Tate–Shafarevich group of $A/K$ is the group
$\Sha(A/K)=\ker\left(H^{1}(G_{K},A)\to\prod_{v}H^{1}(G_{v},A)\right)$
where $v$ ranges over places of $K$. The non-trivial elements of the group $\Sha(A/K)$ parametrise torsors for $A/K$ which have $K_{v}$-rational points for every place $v$, but no $K$-points. In this article we prove the following theorem.
Theorem 1.1.
[…]
Remark 1.2.
The genus $2$ Jacobian $J/\mathbb{Q}$ of conductor $3200^{2}$ in Table 1 was included in an appendix to [KS23] joint with Keller and Stoll where the strong Birch and Swinnerton-Dyer conjecture is also verified for $J$.
$D$ $f(x)$ $\sqrt{N_{J}}$ $E$
$8$ $-10(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 3200 3200.a1
$8$ $165(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 39325 39325.c1
$37$ $-13(27x^{6}-54x^{5}-90x^{4}+228x^{3}+15x^{2}-90x-23)$ 73008 73008.n1
$8$ $-51(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 93925 93925.d1
$8$ $285(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 117325 117325.c1
$8$ $-46(x^{6}+6x^{5}-20x^{4}+240x^{3}+70x^{2}-84x+12)$ 203136 203136.i2
$8$ $-5(3x^{6}+12x^{5}+89x^{4}-56x^{3}-7x^{2}-132x+99)$ 211200 211200.c1
$8$ $465(9x^{6}-12x^{5}+64x^{4}-56x^{3}+136x^{2}-60x+84)$ 216225 432450.ci1
$8$ $-30(11x^{6}-18x^{5}+47x^{4}+6x^{3}+71x^{2}+18x+27)$ 244800 244800.dc1
$8$ $-390(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 270400 270400.dc2
$8$ $-177(3x^{6}+12x^{4}-10x^{3}-12x+11)$ 281961 2819610*
$8$ $-22(x^{6}-24x^{5}+100x^{4}+102x^{3}-80x^{2}-132x-39)$ 302016 302016.p1
$8$ $-6(13x^{6}-116x^{5}-316x^{4}+58x^{3}+264x^{2}-116x+13)$ 313920 313920.bb1
$8$ $-55(13x^{6}-116x^{5}-316x^{4}+58x^{3}+264x^{2}-116x+13)$ 329725 659450*
$8$ $-110(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 387200 8905600*
$8$ $11(x^{6}+6x^{5}+11x^{4}-13x^{2}+6x-2)$ 423984 423984.by1
[…]
The group $\Sha(A/K)$ is a torsion group and is conjectured to be finite. It is conjectured that for every prime number $p>0$ and each integer $g>0$ there exists an absolutely simple abelian variety (i.e., one which is not isogenous over $\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}\mkern-1.5mu}\mkern 1.5mu$ to a product) of dimension $g$ for which $\Sha(A/\mathbb{Q})[p]\neq 0$. Indeed, it is even conjectured that $\Sha(A^{d}/\mathbb{Q})[p]\neq 0$ for a positive proportion of quadratic twists of a fixed abelian variety $A/\mathbb{Q}$ (see e.g., [BKLS21, Conjecture 1.1]).
In spite of this, for general values of $g$ and $p$, constructing an example of an absolutely simple $g$-dimensional abelian variety $A/\mathbb{Q}$ with an $p$-torsion element contained in $\Sha(A/\mathbb{Q})$ is an open problem. By allowing the dimension of $A$ to increase with $p$, Shnidman and Weiss [SW22] construct examples of absolutely simple abelian varieties with $\Sha(A/\mathbb{Q})[p]\neq 0$. Flynn and Shnidman [FS22] extended this result to show $\Sha(A/\mathbb{Q})[p]$ can be arbitrarily large.
When $A$ has dimension $2$, Bruin, Flynn, and Testa [BFT14, Fly15] found examples of absolutely simple genus $2$ Jacobians with $3$ and $5$-torsion in their Tate–Shafarevich groups. Their approach relies on $(p,p)$-descent. That is, for several explicit examples of genus $2$ curves $C/\mathbb{Q}$ they determined the $\psi$-Selmer groups of their Jacobians $J/\mathbb{Q}$ where $\psi$ is a $(p,p)$-isogeny (i.e., a polarised isogeny with kernel isomorphic to $(\mathbb{Z}/p\mathbb{Z})^{2}$) for each $p=3,5$.
However, performing a $(p,p)$-descent becomes computationally costly as $p$ increases, due to the need to perform class and unit group calculations in (a subfield of) the field $\mathbb{Q}(J[\psi])$. Our approach is to instead leverage visibility (see e.g., [CM00, AS02, AS05, Fis16]) to construct absolutely simple genus $2$ Jacobians such that $\Sha(J/\mathbb{Q})[7]\neq 0$.
Let $K$ be a number field and let $A/K$ and $A^{\prime}/K$ be abelian varieties equipped with isogenies $\psi\colon A\to B$ and $\psi^{\prime}\colon A^{\prime}\to B^{\prime}$.
Definition 1.3.
We say that $A/K$ and $A^{\prime}/K$ are $(\psi,\psi^{\prime})$-congruent if there exists a $G_{K}$-equivariant group isomorphism $\phi\colon A[\psi]\to A^{\prime}[\psi^{\prime}]$. We say that $\phi$ is a $(\psi,\psi^{\prime})$-congruence.
In this case let $\Delta=\operatorname{Graph}\phi\subset A\times A^{\prime}$. An element of $\Sha(A/K)$ is said to be visible in the abelian variety $Z=(A\times A^{\prime})/\Delta$ if it is contained in the kernel of the induced homomorphism $H^{1}(G_{K},A)\to H^{1}(G_{K},Z)$.
Visibility is useful for constructing elements of $\Sha(A/K)$ since it allows us to transport information between the Mordell–Weil group of $B^{\prime}/K$ and the Tate–Shafarevich group of $A/K$. More precisely, if $B(K)/\psi A(K)=0$ then, under mild hypotheses applied at the bad primes of $A$ and $A^{\prime}$ and the primes dividing $|\Delta|$ (see [Fis16, Theorem 2.2]), the group $\operatorname{Vis}_{Z}\Sha(A/K)$ of elements of $\Sha(A/K)$ that are visible in $Z=(A\times A^{\prime})/\Delta$ is equal to $B^{\prime}(K)/\psi^{\prime}A^{\prime}(K)$.
The central idea for proving Theorem 1.1 is to construct examples of genus $2$ curves $C/\mathbb{Q}$ with the property that there exists a $(7,\psi)$-congruence between an elliptic curve $E/\mathbb{Q}$ and the Jacobian $J=\operatorname{Jac}(C)$ of $C$, for some isogeny $\psi\colon J\to B$. Assuming that the local conditions are satisfied, it then suffices to show that $B(\mathbb{Q})/\psi J(\mathbb{Q})=0$ and that $E(\mathbb{Q})/7E(\mathbb{Q})\neq 0$ (which in practice is computationally less intensive than performing a $\psi$-descent on $J$). This approach is a mirror to that taken by Fisher [Fis16] who used it to visualise elements of order $7$ in the Tate–Shafarevich groups of elliptic curves. The main technical contribution of this article is construct examples of elliptic curves which are $(7,\psi)$-congruent to a genus $2$ Jacobian.
We ensure the existence of such an isogeny $\psi$ by choosing $J/\mathbb{Q}$ to have real multiplication (RM) by a real quadratic order $\mathcal{O}_{D}$ of fundamental discriminant $D>0$. Suppose that $7$ splits in $\mathcal{O}_{D}$ and we have an embedding $\mathcal{O}_{D}\subset\operatorname{End}_{\mathbb{Q}}(J)$. Writing $(7)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathcal{O}_{D}$ by abuse of notation we write $\mathfrak{p}\colon J\to B$ for the isogeny with kernel consisting of those $P\in J(\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}\mkern-1.5mu}\mkern 1.5mu)$ annihilated by $\mathfrak{p}$. In this case, $\ker\mathfrak{p}$ is isomorphic as a group to $(\mathbb{Z}/7\mathbb{Z})^{2}$ and under suitable hypotheses (see Lemma 2.2) comes equipped with a natural alternating bilinear pairing.
In Section 3.2 we describe an algorithm for determining (a $q$-adic approximation to) a pair of twists of the Klein quartic (the modular curve $X(7)$) which parametrise elliptic curves that are $(7,\mathfrak{p})$-congruent to a fixed genus $2$ Jacobian $J/\mathbb{Q}$ with real multiplication by $\mathcal{O}_{D}$ (our algorithm is subject to the technical hypothesis that $J[\mathfrak{p}]$ is an irreducible $G_{\mathbb{Q}}$-module).
Remark 1.4.
Since abelian surfaces $J/\mathbb{Q}$ with RM by $\mathcal{O}_{D}$ are modular (this follows from Serre’s conjecture [KW09a, KW09b]) we may associate to $J$ a weight $2$ newform with coefficients in $\mathcal{O}_{D}$ and level $\sqrt{N_{J}}$, where $N_{J}$ is the conductor of $J$ (in particular $N_{J}$ is a perfect square).
We compute these twists of $X(7)$ for examples of genus $2$ Jacobians of small conductor provided by the real multiplication families of Bending [Ben98, Ben99] and of Elkies–Kumar [EK14, CFM24]. By searching for rational points on these twists, we find a number of putative examples of $(7,\mathfrak{p})$-congruences between an elliptic curve $E/\mathbb{Q}$ and a genus $2$ Jacobian $J/\mathbb{Q}$. Adapting an approach of Fisher [Fis16, Section 6] we prove these congruences in 5.2.
For an abelian variety $A/\mathbb{Q}$ we write $A^{d}/\mathbb{Q}$ for the quadratic twist of $A$ by a squarefree integer $d\in\mathbb{Z}$. Note that simultaneous quadratic twists of $(7,\mathfrak{p})$-congruent pairs remain $(7,\mathfrak{p})$-congruent (cf. [FK22, Lemma 4.15]). To construct the examples in Theorem 1.1 we simply search for quadratic twists of the examples in 5.2 where there is a rank discrepancy between $E^{d}/\mathbb{Q}$ and $J^{d}/\mathbb{Q}$.
In addition to Theorem 1.1 we prove that there exist examples of such genus $2$ Jacobians with $7$-torsion in their Tate–Shafarevich groups and with real multiplication by $\mathcal{O}_{D}$ for several fundamental discriminants $D>0$.
Theorem 1.5.
For each $D=8$, $29$, $37$, $44$, and $57$ there exists an absolutely simple genus $2$ Jacobian $J/\mathbb{Q}$ with real multiplication by $\mathcal{O}_{D}$ such that $\Sha(J/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$. Examples are furnished by the Jacobians of the curves $C\colon y^{2}=f(x)$ given in Table 2.
$D$ $f(x)$ $\sqrt{N_{J}}$ $E$
$8$ $-10(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 3200 3200.a1
$29$ $-2470(8x^{6}-2x^{5}+68x^{4}+221x^{3}+122x^{2}+986x+1588)$ 40019200 760364800*
$37$ $-39(x^{6}-45x^{4}-68x^{3}+504x^{2}+180x-1193)$ 73008 73008.n1
$44$ $-39(14x^{6}-30x^{5}+85x^{4}+700x^{3}-1325x^{2}+3000x+18000)$ 608400 608400*
$57$ $1479(80x^{6}+279x^{4}+186x^{3}+243x^{2}+324x+108)$ 590609070 7677917910*
Table 2. Examples of genus $2$ curves $C\colon y^{2}=f(x)$ whose Jacobians $J/\mathbb{Q}$ have real multiplication by the quadratic order $\mathcal{O}_{D}$ of discriminant $D$ and such that $\Sha(J/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$. The subgroup of $\Sha(J/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence between $E/\mathbb{Q}$ and $J/\mathbb{Q}$ where $\mathfrak{p}$ divides $7$ in $\mathcal{O}_{D}$. Explicit Weierstrass equations for the corresponding elliptic curves are given in [Fre]. The curves $C/\mathbb{Q}$ were generated from [Ben99, Theorem 4.1] and [EK14, CFM24]. Conductors were computed using [DD19].
In [KS23, A.3] it is observed that the Birch and Swinnerton–Dyer conjecture predicts that $\Sha(J^{-11}/\mathbb{Q})$ contains a subgroup isomoprhic to $(\mathbb{Z}/7\mathbb{Z})^{2}$, where $J/\mathbb{Q}$ is the Jacobian of the genus $2$ curve with LMFDB label 385641.a.385641.1. Since $J$ has RM by $\mathcal{O}_{8}$ it is natural to ask whether the $7$-torsion in $\Sha(J^{-11}/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence with an elliptic curve. By computing the relevant twists of $X(7)$ we give evidence that this is not the case (see Section 6).
Conjecture 1.6.
Let $C/\mathbb{Q}$ be the genus $2$ curve with LMFDB label 385641.a.385641.1 and let $J/\mathbb{Q}$ be its Jacobian. There exists a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$ contained in $\Sha(J^{-11}/\mathbb{Q})$ that is not visible in an abelian $3$-fold.
1.1. Outline of the paper
We begin by discussing several well known facts about the modular curve $X(p)$ in Section 2.1. In Section 2.2, following [PSS07, Section 4.4], we recall the moduli interpretation for twists of $X(p)$, which in Section 2.3 we relate to the torsion of genus $2$ curves with Jacobians having real multiplication (cf. [Fis16]). In Section 2.4 we specialise to the case when $p=7$ and discuss the invariant theory of the Klein quartic $X(7)$ following [Fis14].
In Section 3.2 we present our main algorithm. It takes as input a genus $2$ Jacobian $J/\mathbb{Q}$ with RM by an order in which $7$ splits, and outputs four twists of $X(7)$ which parametrise elliptic curves $(7,\mathfrak{p})$-congruent to $J/\mathbb{Q}$.
The outputs of the algorithm in Section 3.2 are not guaranteed to be correct, however in Section 4 we prove that the output is correct in many cases (for example for the curves in Tables 1 and 2). In particular, in Section 4 we prove that the twists we obtain are isomorphic to those which parametrise elliptic curves $(7,\mathfrak{p})$-congruent to $J$.
In Section 5 we prove Theorems 1.1 and 1.5 by proving that the pairs $(E,J)$ in Tables 1 and 2 are $(7,\mathfrak{p})$-congruent, and by checking that the local hypotheses in [Fis16, Theorem 2.2] are satisfied.
Finally in Section 6 we give explicit examples of the Klein quartic twists for the Jacobian of the genus $2$ curve with LMFDB label 385641.a.385641.1. By searching for rational points on these twists, we give evidence towards 1.6.
1.2. Acknowledgements
This article formed part of my PhD thesis [Fre23, Chapter 6]. I would like to thank my PhD supervisor Tom Fisher for many discussions and for extensive comments on previous versions of this article. I am grateful to Timo Keller for helpful correspondences and to Maria Corte-Real Santos and Ross Paterson for useful comments and discussions.
This work was supported by the Woolf Fisher and Cambridge Trusts, and by the Royal Society through Céline Maistret’s Dorothy Hodgkin Fellowship. The code associated to this article is written in Magma [BCP97] and SageMath [Sage] and is publicly available in the GitHub repository [Fre].
2. The modular curve $X(p)$ and its twists $X_{M}^{\pm}(p)$
We recall a number of standard facts about the modular curve $X(p)$ and its twists $X_{M}^{\pm}(p)$ following e.g., [PSS07] and [Fis14]. In the case when $p=7$ the curve $X(7)$ is the Klein quartic [Kle78] (see [Elk99] for a detailed discussion).
[…]
2.1. The modular curve $X(p)$
Let $E/K$ be an elliptic curve defined over a field $K$ of characteristic zero. If $p$ is a prime number, we equip $E[p]$ with the structure of a symplectic abelian group via the $p$-Weil pairing $e_{E,p}\colon E[p]\times E[p]\to\mu_{p}$.
Let $Y(p)/\mathbb{Q}$ denote the geometrically irreducible (non-compact) modular curve parametrising elliptic curves with full (symplectic) level $p$ structure. Explicitly, for each field $K/\mathbb{Q}$ the $K$-points on $Y(p)$ parametrise (isomorphism classes of) pairs $(E,\iota)$ where $E/K$ is an elliptic curve and $\iota\colon\mu_{p}\times\mathbb{Z}/p\mathbb{Z}\cong E[p]$ is a $G_{K}$-equivariant isomorphism of symplectic abelian groups. Let $X(p)$ denote the smooth compactification of $Y(p)$.
The group $\Gamma\!_{p}$ of symplectic automorphisms of $\mu_{p}\times\mathbb{Z}/p\mathbb{Z}$ acts naturally on $Y(p)$ on the right via $(E,\iota)\mapsto(E,\gamma\iota)$. As an abstract group $\Gamma\!_{p}$ is isomorphic to $\operatorname{SL}_{2}(\mathbb{Z}/p\mathbb{Z})$, but it comes equipped with a non-trivial action of $\operatorname{Gal}(\mathbb{Q}(\mu_{p})/\mathbb{Q})$. The matrix ±$\pm I$ acts trivially on $Y(p)$ and therefore the action of $\Gamma\!_{p}$ factors through $\Gamma\!_{p}/\{\pm I\}$. This action extends to an action on $X(p)$, and the quotient realises the forgetful morphism $X(p)\to X(1)$ given by taking $j$-invariants.
2.2. The twist $X_{M}^{\pm}(p)$
Let $(M,e_{M})$ be a symplectic abelian group over $K$. Let $r$ be an integer coprime to $p$ and suppose that there exists a $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-isomorphism $\phi\colon M\cong\mu_{p}\times\mathbb{Z}/p\mathbb{Z}$ such that $\langle\phi(P),\phi(Q)\rangle=e_{M}(P,Q)^{r}$ for each $P,Q\in\mu_{p}\times\mathbb{Z}/p\mathbb{Z}$.
By the twisting principle we may attach to $\phi$ a cohomology class $\xi\in H^{1}(G_{K},\Gamma\!_{p})$. We have an inclusion $\Gamma_{p}/\{\pm I\}\hookrightarrow\operatorname{Aut}(X(p))$ and therefore an induced map on cohomology $H^{1}(G_{K},\Gamma\!_{p})\to H^{1}(G_{K},\operatorname{Aut}(X(p)))$. The image of $\xi$ corresponds (again by the twisting principle) to a twist $X_{M}^{r}(p)$ of $X(p)$.
The following lemma is well known and follows by construction (cf. [PSS07, Section 4.4]).
Lemma 2.1.
For each field $L/K$ the $L$-rational points on $X_{M}^{r}(p)$ correspond to pairs $(E,\phi)$ where $E/L$ is an elliptic curve and $\phi\colon M\cong E[p]$ is an isomorphism of $G_{L}$-modules for which $e_{E,p}(\phi(P),\phi(Q))=e_{M}(P,Q)^{r}$ for each $P,Q\in M$.
If $a$ is an integer coprime to $p$, pre-composing an isomorphism $\phi\colon M\cong E[p]$ with the multiplication-by-$a$-map on $M$ yields an isomorphism $\phi^{\prime}$ for which $e_{E,p}(\phi^{\prime}(P),\phi^{\prime}(Q))=e_{M}(P,Q)^{a^{2}r}$. It therefore suffices to consider the class of $r$ in $(\mathbb{Z}/p\mathbb{Z})^{\times}$ modulo squares. We write $X_{M}(p)=X_{M}^{+}(p)$ when $r$ is a square in $(\mathbb{Z}/p\mathbb{Z})^{\times}$ and $X_{M}^{-}(p)$ when $r$ is not a square in $(\mathbb{Z}/p\mathbb{Z})^{\times}$.
2.3. The twist $X^{r}_{J[\mathfrak{p}]}(p)$
Let $C/K$ be a genus $2$ curve and let $J=\operatorname{Jac}(C)$ be the Jacobian of $C$. Let $\widehat{J}$ denote the dual of $J$ and equip $J$ with the canonical principal polarisation $\lambda\colon J\to\widehat{J}$ arising from the theta divisor. The principal polarisation $\lambda$ induces an involution on the endomorphism ring of $J$ known as the Rosati involution. Precisely, if $\psi\in\operatorname{End}(J)$ then the Rosati involution is given by $\psi\mapsto\psi^{\dagger}=\lambda^{-1}\hat{\psi}\lambda$. Here $\hat{\psi}\colon\widehat{J}\to\widehat{J}$ denotes the dual isogeny.
Let $D\equiv 0,1\pmod{4}$ be a positive non-square integer and let $\mathcal{O}_{D}$ be the quadratic ring of discriminant $D$. We say that $J$ has real multiplication (RM) by $\mathcal{O}_{D}$ if there exists an inclusion $\mathcal{O}_{D}\hookrightarrow\operatorname{End}^{\dagger}_{K}(J)$ where $\operatorname{End}_{K}^{\dagger}(J)\subset\operatorname{End}_{K}(J)$ is the subring of endomorphisms fixed by the Rosati involution.
The choice of principal polarisation $\lambda$ induces the alternating, bilinear $p$-Weil pairing $e_{J,p}\colon J[p]\times J[p]\to\mu_{p}$.
Lemma 2.2.
Let $J/K$ be a genus $2$ Jacobian with RM by $\mathcal{O}_{D}$. Suppose that $p$ is a prime number such that $p$ splits as a product $(p)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathcal{O}_{D}$. Then there exists an isomorphism of $G_{K}$-modules $J[p]\cong J[\mathfrak{p}]\oplus J[\bar{\mathfrak{p}}]$. Moreover, the $p$-Weil pairing $e_{J,p}$ restricts to an alternating pairing $J[\mathfrak{p}]\times J[\mathfrak{p}]\to\mu_{p}$, and likewise for $J[\bar{\mathfrak{p}}]$.
Proof.
This is [CFM24, Lemma 3.4], cf. the proof of [Fis16, Proposition 6.1] when $\mathcal{O}_{D}=\mathbb{Z}[\sqrt{2}]$. ∎
Note that by Lemma 2.2 we may define the twists $X_{J[\mathfrak{p}]}^{\pm}(p)$ and $X_{J[\bar{\mathfrak{p}}]}^{\pm}(p)$ which (by Lemma 2.1) parametrise elliptic curves which are $(p,\mathfrak{p})$-congruent (respectively $(p,\bar{\mathfrak{p}})$-congruent) to the genus $2$ Jacobian $J/K$.
2.4. Explicit twisting for $X(7)$
[…]
$\frac{1}{\sqrt{-7}}\begin{pmatrix}\zeta_{7}-\zeta_{7}^{6}&\zeta_{7}^{2}-\zeta_{7}^{5}&\zeta_{7}^{4}-\zeta_{7}^{3}\\ \zeta_{7}^{2}-\zeta_{7}^{5}&\zeta_{7}^{4}-\zeta_{7}^{3}&\zeta_{7}-\zeta_{7}^{6}\\ \zeta_{7}^{4}-\zeta_{7}^{3}&\zeta_{7}-\zeta_{7}^{6}&\zeta_{7}^{2}-\zeta_{7}^{5}\\ \end{pmatrix}\quad\text{and}\quad\begin{pmatrix}\zeta_{7}&0&0\\ 0&\zeta_{7}^{4}&0\\ 0&0&\zeta_{7}^{2}\\ \end{pmatrix}.$
Composing with the natural $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-isomorphism $\Gamma\!_{7}\cong\operatorname{SL}_{2}(\mathbb{Z}/7\mathbb{Z})$ gives a $G_{K}$-equivariant homomorphism $\rho\colon\Gamma\!_{7}\to\operatorname{GL}_{2}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu)$. When $\Gamma\!_{7}$ acts on $\mathbb{P}^{2}$ via ${\rho}$ the Klein quartic curve $X(7)\subset\mathbb{P}^{2}$ given by
$X(7):x_{0}^{3}x_{1}+x_{1}^{3}x_{2}+x_{0}x_{2}^{3}=0$
is fixed.
Let $\phi\colon M\cong\mu_{7}\times\mathbb{Z}/7\mathbb{Z}$ be a symplectic $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-isomorphism. Let $\xi\colon G_{K}\to\operatorname{GL}_{3}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu)$ denote the cocycle obtained from $\phi$ by the twisting principle. Explicitly, $\xi$ may be taken to be the cocycle $\sigma\mapsto{\rho}(\sigma(\phi)\phi^{-1})$ where $\sigma(\phi)\colon M\cong\mu_{7}\times\mathbb{Z}/7\mathbb{Z}$ is the $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-isomorphism given by $P\mapsto\sigma(\phi(\sigma^{-1}P))$. By Hilbert’s Theorem 90 the cocycle $\xi$ is a coboundary, that is, there exists a matrix $A\in\operatorname{GL}_{3}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu)$ such that $\sigma(A^{-1})A=\xi(\sigma)$ for every $\sigma\in G_{K}$.
Lemma 2.3.
The curves $X_{M}(7)$ and $X_{M}^{-}(7)$ are isomorphic to the images of the morphisms $X(7)\to\mathbb{P}^{2}$ given by
$\mathbf{x}\mapsto A\mathbf{x}\quad\text{and}\quad\mathbf{x}\mapsto A^{-{T}}\mathbf{x}$
respectively. Here $\mathbf{x}$ is a point in $\mathbb{P}^{2}$ written as a column vector and $A^{-{T}}$ denotes the inverse transpose of $A$.
Proof.
The proof is identical to [Fis14, Lemma 3.2]. ∎
Finally, we note that it is simple to recover the moduli interpretation for a given twist of $X(7)$ following [Fis14, Section 4.1] (cf. [PSS07, Section 7.1]).
Lemma 2.4.
Let $\mathcal{X}/K$ be a twist of the Klein quartic given by the vanishing of a homogeneous quartic polynomial $\mathcal{F}(x_{0},x_{1},x_{2})$ in $\mathbb{P}^{2}$. The $j$-invariant map $\mathcal{X}\to X(1)$ is defined over $K$ and is given by $1728\frac{c_{4}^{3}}{c_{4}^{3}-c_{6}^{2}}$ where $c_{4}$ and $c_{6}$ are defined by
$D=\frac{-1}{54}\times\left|\begin{matrix}\frac{\partial^{2}\mathcal{F}}{\partial x_{0}^{2}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{0}\partial x_{1}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{0}\partial x_{2}}\\[10.00002pt] \frac{\partial^{2}\mathcal{F}}{\partial x_{0}x_{1}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{1}^{2}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{1}\partial x_{2}}\\[10.00002pt] \frac{\partial^{2}\mathcal{F}}{\partial x_{0}x_{2}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{1}\partial x_{2}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{2}^{2}}\end{matrix}\right|,\quad c_{4}=\frac{1}{9}\times\left|\begin{matrix}\frac{\partial^{2}\mathcal{F}}{\partial x_{0}^{2}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{0}\partial x_{1}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{0}\partial x_{2}}&\frac{\partial D}{\partial x_{0}}\\[10.00002pt] \frac{\partial^{2}\mathcal{F}}{\partial x_{0}\partial x_{1}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{1}^{2}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{1}\partial x_{2}}&\frac{\partial D}{\partial x_{1}}\\[10.00002pt] \frac{\partial^{2}\mathcal{F}}{\partial a\partial x_{2}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{1}\partial x_{2}}&\frac{\partial^{2}\mathcal{F}}{\partial x_{2}^{2}}&\frac{\partial D}{\partial x_{2}}\\[10.00002pt] \frac{\partial D}{\partial x_{0}}&\frac{\partial D}{\partial x_{1}}&\frac{\partial D}{\partial x_{2}}&0\end{matrix}\right|,\quad c_{6}=\frac{1}{14}\times\left|\begin{matrix}\frac{\partial\mathcal{F}}{\partial x_{0}}&\frac{\partial\mathcal{F}}{\partial x_{1}}&\frac{\partial\mathcal{F}}{\partial x_{2}}\\[10.00002pt] \frac{\partial D}{\partial x_{0}}&\frac{\partial D}{\partial x_{1}}&\frac{\partial D}{\partial x_{2}}\\[10.00002pt] \frac{\partial c_{4}}{\partial x_{0}}&\frac{\partial c_{4}}{\partial x_{1}}&\frac{\partial c_{4}}{\partial x_{2}}\end{matrix}\right|.$
3. Computing approximations to twists of $X(7)$
Fix a primitive $7^{\text{th}}$ root of unity $\zeta_{7}$, and let $J/\mathbb{Q}$ be the Jacobian of a genus $2$ curve $C/\mathbb{Q}$. Suppose that $J$ has RM by the quadratic order $\mathcal{O}_{D}$ of fundamental discriminant $D>0$. Suppose that $7$ splits in $\mathcal{O}_{D}$ and that we have a factorisation $(7)=\mathfrak{p}\bar{\mathfrak{p}}$ (where $\bar{\mathfrak{p}}$ denotes the conjugate of $\mathfrak{p}$). We assume throughout this section that $J[\mathfrak{p}]$ is an irreducible $G_{\mathbb{Q}}$-module.
Let $\mathcal{K}=J/\{\pm 1\}$ denote the Kummer surface of $J$. We identify $\mathcal{K}$ with a singular quartic surface in $\mathbb{P}^{3}$ by the embedding in [CF96, (3.1.8)]. Let $x_{J}$ denote the quotient morphism $J\to\mathcal{K}$. If $\psi\colon J\to J^{\prime}$ is an isogeny we write $\mathcal{K}[\psi]=x_{J}(J[\psi])$.
3.1. Hilbert’s Theorem 90 is effective
To compute twists of $X(7)$ using Lemma 2.3 we will need to compute matrices which realise a given cocycle as a coboundary. Towards this, note that the standard proof of Hilbert’s Theorem 90 is “nearly” effective. Indeed following [Ser79, Proposition X.3] let $L/K$ be a finite extension of infinite fields, let $\xi\in H^{1}(\operatorname{Gal}(L/K),\operatorname{GL}_{n}(L))$ be a $1$-cocycle, and choose an element $c\in\operatorname{GL}_{n}(L)$. Define a matrix $b\in\mathrm{M}_{n}(L)$ by the Poincaré series
$b=\sum_{\sigma\in\operatorname{Gal}(L/K)}\xi(\sigma)\sigma(c).$
If $c$ is chosen so that $b$ is invertible, then $\xi(\sigma)=\sigma(b)^{-1}b$ and it is immediate that $\xi$ is a coboundary. If $L$ is infinite the linear independence of field embeddings guarantees the existence of $c$. Indeed, for a fixed cocycle $\xi$, the failure of $b$ to be invertible is a Zariski closed condition on the matrix $c$. In particular, for a generic choice of $c$, the matrix $b$ will be invertible. The proof therefore suggests an algorithm.
Algorithm 3.1 Generating matrices which realise Hilbert’s Theorem 90.
Input: A cocycle $\xi\in H^{1}(\operatorname{Gal}(L/K),\operatorname{GL}_{n}(L))$.
Output: A matrix $b\in\operatorname{GL}_{n}(L)$ such that $\xi(\sigma)=\sigma(b)^{-1}b$ for each $\sigma\in\operatorname{Gal}(L/K)$.
1:   Choose “randomly” a matrix $c\in\operatorname{GL}_{n}(L)$.
2:  Compute the matrix $b=\sum_{\sigma}\xi(\sigma)\sigma(c)$.
3:  if $\det(b)\neq 0$ then
4:     return  $b$
5:  else
6:     Return to step 1.
7:  end if
Remark 3.2.
In step 1 of Algorithm 3.1 the user must choose a matrix $c\in\operatorname{GL}_{n}(L)$. In our application the extension $L/\mathbb{Q}$ will be a finite extension and we will have access to a LLL-reduced $\mathbb{Z}$-basis $\{a_{1},...,a_{m}\}$ for the ring of integers $\mathcal{O}_{L}\subset L$. We choose “small” elements of $L$ by generating a tuple $x_{1},...,x_{m}\in\{0,\pm 1\}$ and considering the element $\sum_{i=1}^{m}x_{i}a_{i}\in\mathcal{O}_{L}$. This approach extends to choosing a matrix $c\in\operatorname{GL}_{n}(L)$ by choosing the $n^{2}$ entries as described (in practice, we also choose almost all $x_{i}$ to be equal to zero).
Remark 3.3.
In principle the recursive step in Algorithm 3.1 may be called many times. In practice, however, we have found very few instances when Algorithm 3.1 fails to terminate in one iteration.
3.2. The main algorithm
We write $L=\mathbb{Q}(\mathcal{K}[\mathfrak{p}])$. Note that for generic $J/\mathbb{Q}$, we have an isomorphism of abstract groups $\operatorname{Gal}(L/\mathbb{Q})\cong\operatorname{GL}_{2}(\mathbb{Z}/7\mathbb{Z})/\{\pm 1\}$. Suppose that we have degree $24$ polynomials $g_{1}(t)$, $g_{2}(t)$, $g_{3}(t)\in\mathbb{Q}[t]$ such that
(1) $\mathcal{K}[\mathfrak{p}]\subset\{(1:\alpha_{1}:\alpha_{2}:\alpha_{3})\in\mathcal{K}:g_{1}(\alpha_{1})=g_{2}(\alpha_{2})=g_{3}(\alpha_{3})=0\}\cup\{(0:0:0:1)\}.$
In Section 3.3 we discuss how the polynomials $g_{i}(t)$ may be computed.
We fix an auxiliary prime $q\neq 7$ not dividing the discriminant of $g_{1}(t)$ and at which $J$ has good reduction. Further suppose that the minimal polynomial of $\zeta_{7}$ is irreducible over $\mathbb{F}_{q}$ and that $L$ is equal to the splitting field of $g_{1}(t)$. Let $\mathfrak{q}$ be a prime of $L$ dividing $q$ and denote by $L_{\mathfrak{q}}$ and $l_{\mathfrak{q}}$ the completion and residue field of $L$ at $\mathfrak{q}$ respectively.
Our algorithm proceeds as follows:
  1. 1:
    We compute the Galois group of $g_{1}(t)$ using GaloisGroup in Magma. This gives a group $G\subset S_{24}$ and a $G$-set $\{r_{1},...,r_{24}\}$ of ($\mathfrak{q}$-adic approximations to) the roots of $g_{1}(t)$ in $L_{\mathfrak{q}}$ such that $G$ gives the action of $\operatorname{Gal}(L/\mathbb{Q})$ on the roots of $g_{1}(t)$ in $L$.
  2. 2:
    We compute (a $\mathfrak{q}$-adic approximation to) an embedding $\mathbb{Q}_{q}(\zeta_{7})\hookrightarrow L_{\mathfrak{q}}$ by computing a $\mathfrak{q}$-adic approximation to $\zeta_{7}$.
  3. 3:
    We compute a $\mathbb{Z}/7\mathbb{Z}$-basis $\{\bar{P},\bar{Q}\}$ for $J(l_{\mathfrak{q}})[\mathfrak{p}]$ such that $e_{J,7}(\bar{P},\bar{Q})=\zeta_{7}$ (via the Magma intrinsic WeilPairing). This uniquely determines a pair $x_{J}(P),x_{J}(Q)\in\mathcal{K}(L_{\mathfrak{q}})$ which reduce modulo $\mathfrak{q}$ to $x_{J}(\bar{P})$ and $x_{J}(\bar{Q})$ respectively.
  4. 4:
    Let $\phi\colon J[\mathfrak{p}]\cong\mu_{7}\times\mathbb{Z}/7\mathbb{Z}$ be the isomorphism given by $P\mapsto(\zeta_{7},0)$ and $Q\mapsto(1,1)$. We explicitly determine the 1-cocycle
    $\xi\colon\operatorname{Gal}(L/\mathbb{Q})\cong G\to\operatorname{GL}_{3}(\mathbb{Q}(\zeta_{7}))$
    given by $\sigma\mapsto\rho(\sigma(\phi)\phi^{-1})$ where $\rho$ is the representation given in Section 2.4.
  5. 5:
    We compute (a $\mathfrak{q}$-adic approximation to) a matrix $A\in\operatorname{GL}_{3}(L)\subset\operatorname{GL}_{3}(L_{\mathfrak{q}})$ which realises $\xi$ as a coboundary, using Algorithm 3.1.
  6. 6:
    We twist $X(7)$ by $A$ and $A^{-T}$ to obtain curves $\mathcal{X}^{\pm}\subset\mathbb{P}^{2}_{L_{\mathfrak{q}}}$. By Lemma 2.3 these twists are $\mathfrak{q}$-adic approximations to the twists $X_{J[\mathfrak{p}]}^{\pm}(7)\subset\mathbb{P}^{2}_{\mathbb{Q}}$. After normalising each equation so that the first non-zero coefficient is equal to $1$, the curves obtained therefore have coefficients in $\mathbb{Q}$ (up to a numerical error).
  7. 7:
    We recognise the coefficients of the twists $\mathcal{X}^{\pm}$ as rational numbers using the LLL algorithm.
  8. 8:
    We minimise and reduce the models for $\mathcal{X}^{\pm}$ using the algorithm of Elsenhans–Stoll [ES24], which is implemented in Magma as MinRedTernaryForm.
Remark 3.4.
Computing the matrix $A\in\operatorname{GL}_{3}(L_{\mathfrak{q}})$ must be done with some care in order to control coefficient explosion (and to minimise the $\mathfrak{q}$-adic precision we must carry throughout the calculation). In our implementation we assume that $J[\mathfrak{p}]$ is an irreducible $G_{\mathbb{Q}}$-module and (heavily) rely on the following observation:
Let $\alpha$ be a root of $g_{1}(t)$ (so that $L$ is the splitting field of $\mathbb{Q}(\alpha)$). Since the Galois module $J[\mathfrak{p}]$ is irreducible, there exists a subfield $\mathbb{Q}\subset K\subset\mathbb{Q}(\alpha)$ (which is unique up to conjugacy) such that $\mathcal{K}[\mathfrak{p}]$ contains a $\operatorname{Gal}(L/K)$-stable “cyclic subgroup” (or more precisely, the image of a cyclic subgroup of $J[\mathfrak{p}]$). From the properties of the Weil pairing we have $L=\widetilde{K}(\zeta_{7})$, where $\widetilde{K}$ is the Galois closure of $K$ in $L$. We first compute an LLL-reduced basis $\{1,k_{1},...,k_{7}\}$ for $K/\mathbb{Q}$. The elements $k\zeta_{7}^{i}$ span $L/\mathbb{Q}$ where $0\leq i\leq 6$ and $k$ ranges over the $\operatorname{Gal}(L/\mathbb{Q})$-conjugates of $k_{j}$ for each $1\leq j\leq 7$. We then apply Algorithm 3.1 noting Remark 3.2.
3.3. Computing the polynomials $g_{i}(t)$
It remains to describe how the polynomials $g_{1}(t)$, $g_{2}(t)$, and $g_{3}(t)$ which cut out $\mathcal{K}[\mathfrak{p}]$ may be computed. Let $C/\mathbb{Q}$ be a genus $2$ curve given by a Weierstrass equation $C:y^{2}=f(x)$ whose Jacobian $J/\mathbb{Q}$ has RM by $\mathcal{O}_{D}$, and such that $7$ splits in $\mathcal{O}_{D}$.
[…]
The model for the Kummer surface $\mathcal{K}$ of $J$ given in [CF96, Chapter 3] and maps $\mathfrak{D}$ to the point $(1:x_{1}+x_{2}:x_{1}x_{2}:\beta_{0})\in\mathcal{K}$ where $\beta_{0}\in\mathbb{Q}(x_{1},x_{2},y_{1},y_{2})$ is the rational function in [CF96, (3.1.4)]. Using the LLL algorithm we compute polynomials $h_{1}(t)$, $h_{2}(t)$, and $h_{3}(t)\in\mathbb{Q}(t)$ which approximate the minimal polynomials of $x_{1}+x_{2}$, $x_{1}x_{2}$, and $\beta_{0}$ (in particular we utilise the Magma function MinimalPolynomial). Using the description of the multiplication-by-$n$-map on $\mathcal{K}$ given in [CF96, Chapter 3] it is simple to verify (unconditionally) that the polynomials $h_{i}(t)$ cut out a $7$-torsion point in $\mathcal{K}(\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}\mkern-1.5mu}\mkern 1.5mu)$. Polynomials $g_{i}(t)$ which satisfy (1) are then the product over the distinct polynomials $h_{i}(t)$ occurring for such divisors $\mathfrak{D}$.
When $D=8$ we also have the following approach which avoids the numerical instability issues which can occur when using AnalyticJacobian.
3.3.1. A numerically stable approach when $D=8$
Fix an isomorphism $\mathcal{O}_{8}\cong\mathbb{Z}[\sqrt{2}]$. The prime number $7$ is a norm from $\mathbb{Z}[\sqrt{2}]$ and we may write $(7)=(3+\sqrt{2})(3-\sqrt{2})$. Let $[\sqrt{2}]\colon J\to J$ denote the multiplication-by-$\sqrt{2}$-map on $J$. The morphism $[\sqrt{2}]$ is a Richelot isogeny and using the approach in [Nic18, Section 5.7] (which is implemented in [Nic]) we determine explicit polynomials giving the morphism $[\sqrt{2}]\colon\mathcal{K}\to\mathcal{K}$ induced by the action of $\sqrt{2}$ on $J$.
Remark 3.5.
By interpolation it is not difficult to give an explicit morphism $\mathcal{K}\to\mathcal{K}$ realising the $\sqrt{2}$-action on the Jacobian of the generic member of the generic family of genus $2$ curves $\mathcal{C}/\mathbb{Q}(A,P,Q)$ provided by Bending [Ben98, Ben99]. We record explicit equations for this (generic) morphism in [Fre].
Formulae for the multiplication-by-$3$-map $[3]\colon\mathcal{K}\to\mathcal{K}$ are given in [CF96, Section 3.5]. Note that $\mathcal{K}[3+\sqrt{2}]\cup\mathcal{K}[3-\sqrt{2}]$ is exactly the set $\{P\in\mathcal{K}:3P=\sqrt{2}P\}$. By taking successive resultants (and fixing a choice of sign so that $\mathfrak{p}=(3\pm\sqrt{2})$) it is simple to compute polynomials $g_{1}(t),g_{2}(t),g_{3}(t)\in\mathbb{Q}[t]$ satisfying (1).
3.4. Outputs of the main algorithm
We provide a Magma implementation of the algorithm described in Section 3.2. The main non-trivial input in the algorithm is a genus $2$ curve $C/\mathbb{Q}$ with RM by an order $\mathcal{O}_{D}$ in which $7$ splits. The fundamental discriminants $D<100$ for which this occurs are $D=8$, $29$, $37$, $44$, $53$, $57$, $60$, $65$, $85$, $88$, $92$, and $93$.
[…]
We run the algorithm in Section 3.2 for each curve recorded in [Fre].
Remark 3.6.
It would be interesting to compute the twists $X_{\mathcal{J}[\mathfrak{p}]}^{\pm}(7)$ for the Jacobian $\mathcal{J}$ of the generic curves $\mathcal{C}/\mathbb{Q}(a,b,c)$ with RM by $\mathcal{O}_{D}$ given in [Ben98, Ben99] and [CFM24] for each $D=8$, $29$, $37$, $44$, and $53$ (i.e., those $D$ where $7$ splits in $\mathcal{O}_{D}$ and for which [CFM24] gives a generic model for a curve by $\mathcal{O}_{D}$). Unfortunately the algorithm we describe is ill suited to this task. One might hope to interpolate over twists computed for a large number of specialisations. However these twists are only defined up to the action of $\operatorname{Aut}_{\mathbb{Q}}(\mathbb{P}^{2})\cong\operatorname{PGL}_{3}(\mathbb{Q})$ and our algorithm for generating matrices which satisfy Hilbert’s Theorem 90 does not do so in a compatible way (it requires a choice of $\mathfrak{p}|7$, a choice of basis for $J[\mathfrak{p}]$, and a “randomly” generated matrix).
4. Proving twists of $X(7)$ are isomorphic to $X_{M}^{\pm}(7)$
Let $M/\mathbb{Q}$ be an irreducible $G_{\mathbb{Q}}$-module and let $\mathcal{X}/\mathbb{Q}$ be a plane quartic curve (in our case we will take $M=J[\mathfrak{p}]$ and $\mathcal{X}$ to be an output of the algorithm in Section 3.2). We now outline an approach for proving that $\mathcal{X}$ is isomorphic to a twist $X_{M}^{\pm}(7)$ of the Klein quartic (for some choice of sign). We assume that $\mathcal{X}$ is a twist of $X(7)$ (note that this is simple to check by computing Dixmier–Ohno invariants [Dix87, Ohn, Els15] in Magma).
For the purpose of proving Theorem 1.1 it suffices to consider only the case when $\mathcal{X}$ has a rational point (i.e., it suffices to recall [Fis16, Lemma 6.2], see Lemma 4.2 below). In Section 4.2 we note how one may prove that a twist is isomorphic to $X_{M}^{\pm}(7)$ more generally.
4.1. When $\mathcal{X}$ has a rational point
Let $K$ be a field of characteristic zero. Suppose that $\mathcal{X}$ has a $K$-rational point which corresponds (through the moduli interpretation in Lemma 2.4) to an elliptic curve $E/K$ (defined up to quadratic twist) with $j$-invariant $j(E)\neq 0,1728,\infty$. In this case, the following lemma reduces the problem of showing that $\mathcal{X}$ is isomorphic to $X_{M}^{\pm}(7)$ to the problem of showing that $X_{E}(7)$ is isomorphic to $X_{M}^{\pm}(7)$.
Lemma 4.1.
Let $\mathcal{X}/K$ be a twist of $X(7)$ and suppose that there exsists a point $P\in\mathcal{X}(K)$ with $j(P)\neq 0,1728,\infty$. If $E/K$ is an elliptic curve with $j(E)=j(P)$ then $\mathcal{X}$ is isomorphic to $X_{E}(7)$ over $K$.
Proof.
Let $\varphi\colon\mathcal{X}\cong X_{E}(7)$ be a $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-isomorphism. By composing with a $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-automorphism of $X_{E}(7)$ we may assume that $\varphi(P)$ is equal to the tautological point $Q=(E,\mathrm{id})\in X_{E}(7)(K)$. Since $P$ and $Q$ are $K$-rational, for each $\sigma\in G_{K}$ we have $\sigma\varphi\sigma^{-1}(P)=Q$, so that $\varphi^{-1}\sigma\varphi\sigma^{-1}(P)=P$ for each $\sigma\in G_{K}$. Since $j(P)\neq 0,1728,\infty$ the only $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-automorphism of $\mathcal{X}$ which fixes $P$ is the identity. In particular, $\varphi=\sigma\varphi\sigma^{-1}$ for all $\sigma\in G_{K}$ and therefore $\varphi$ is defined over $K$. ∎
To show that $X_{E}(7)$ is isomorphic to $X_{M}^{\pm}(7)$ for some choice of sign, it suffices to show that $E[7]$ is isomorphic to $M$ as a $G_{\mathbb{Q}}$-module, up to quadratic twist. We recall the following lemma of Fisher (based on an argument of Serre using Goursat’s lemma [Ser72, Lemme 8]) which allows us to prove such congruences, up to quadratic twists.
Lemma 4.2.
Let $K$ be a number field and let $M$ be a $G_{K}$-module which is isomorphic as an abstract group to $(\mathbb{Z}/p\mathbb{Z})^{2}$ for some $p\geq 5$. Suppose that $M$ comes equipped with a ($G_{K}$-equivariant) alternating pairing $M\times M\to\mu_{p}$. Let $E/K$ be an elliptic curve with surjective mod $p$ Galois representation, let $x_{M}\colon M\to M/\{\pm 1\}$, and let $x_{E}\colon E\to\mathbb{P}^{1}$ be the quotient by {±$\{\pm 1\}$. If there exist non-identity elements $P\in M$ and $Q\in E[p]$ such that $K(x_{M}(P))=K(x_{E}(Q))$ then there exists a quadratic twist $E^{d}$ of $E$ such that $M\cong E^{d}[p]$.
Proof.
This follows immediately from [Fis16, Lemma 6.2] (cf. [Fis16, Proposition 6.1]). Note that the hypothesis that $K=\mathbb{Q}$ in [Fis16, Lemma 6.2] is not used. ∎
4.2. When $\mathcal{X}$ has no rational points
We rely on the approach in Section 4.1 together with the following criterion. In practice when $X$ and $Y$ are twists of $X(7)$ defined over $\mathbb{Q}$ it is simple to find number fields for which the statement holds. In this case there exist infinitely many points on $X$ and $Y$ defined over quartic fields. One expects that if $X$ and $Y$ are isomorphic (and have no non-trivial automorphisms defined over $\mathbb{Q}$), then for a generic such field the conditions of the lemma hold.
Lemma 4.3.
Let $X/K$ and $Y/K$ be (geometrically integral) curves defined over a number field $K$. Suppose that there exist extensions $L_{1},L_{2}/K$ for which $L_{1}\cap L_{2}=K$ and such that we have isomorphisms $\varphi_{i}\colon X_{L_{i}}\cong Y_{L_{i}}$ for each $i=1,2$. If $X$ (or $Y$) does not admit a non-trivial automorphism over the compositum $L_{1}L_{2}$, then $X$ and $Y$ are isomorphic over $K$.
Proof.
The assumption on the automorphism group of $X$ over $L_{1}L_{2}$ implies that the composition $\varphi_{1}^{-1}\varphi_{2}$ is the identity, and therefore over $L_{1}L_{2}$ we have an equality $\varphi_{1}=\varphi_{2}$. But then $\varphi_{1}$ is defined over $L_{1}\cap L_{2}=K$ and the claim follows. ∎
Proposition 4.4.
Conider any of the data in [Fre, data/twists.m] which consists of
  1. ()
    a genus $2$ curve $C/\mathbb{Q}$,
  2. ()
    a fundamental discriminant $D>0$ such that the Jacobian of $C$ has RM by $\mathcal{O}_{D}$, and
  3. ()
    a twist $\mathcal{X}/\mathbb{Q}$ of the Klein quartic.
Then $\mathcal{X}$ is isomorphic over $\mathbb{Q}$ to $X_{J[\mathfrak{p}]}^{\pm}(7)$ for some choice of sign and choice of prime $\mathfrak{p}\subset\mathcal{O}_{D}$ above $7$.
Proof.
If $\mathcal{X}$ has a $\mathbb{Q}$-rational point of small height corresponding to an elliptic curve $E/\mathbb{Q}$, we apply Lemma 4.2. In each case, applying [Ser72, Proposition 19] at several good primes (or using Zywina’s algorithm [Zyw22]) suffices to show that the mod $7$ Galois representation attached to $E/\mathbb{Q}$ is surjective. In the electronic data we exhibit an explicit isomorphism between the fields $x_{J}(P)$ and $x_{E}(Q)$ for some $P\in J[\mathfrak{p}]$ and $Q\in E[7]$ (note that a minimal polynomial for the extension $\mathbb{Q}(x_{J}(P))/\mathbb{Q}$ was computed in the course of the algorithm in Section 3.3).
The general case proceeds similarly. Taking hyperplane sections of $\mathcal{X}$ we construct non-isomorphic quartic fields $L_{1},L_{2}/\mathbb{Q}$ over which $\mathcal{X}$ obtains a point and such that $L_{i}$ contains no non-trivial subfield for each $i=1,2$ (in particular $L_{1}\cap L_{2}=\mathbb{Q}$ and $L_{1}\cap\mathbb{Q}(\zeta_{7})=L_{2}\cap\mathbb{Q}(\zeta_{7})=\mathbb{Q}$). These points correspond to elliptic curves $E_{1}/L_{1}$ and $E_{2}/L_{2}$ whose mod $7$ Galois representations may be seen to be surjective by applying [Ser72, Proposition 19] at several places of good reduction. Applying Lemma 4.2 as above shows that $E_{1}$ and $E_{2}$ are $(7,\mathfrak{p})$-congruent to $J$, up to a quadratic twist. Since the mod $7$ Galois representations of $E_{1}/L_{1}$ and $E_{2}/L_{2}$ are surjective we have $\mathbb{Q}(\mathcal{K}[\mathfrak{p}])\cap L_{1}L_{2}=\mathbb{Q}$, where $\mathcal{K}=J/\{\pm 1\}$ is the Kummer surface of $J$.
It follows from the construction that $\operatorname{Aut}(X_{J[\mathfrak{p}]}^{r}(7))$ is isomorphic (as a $G_{\mathbb{Q}}$-module) to the group $\operatorname{Aut}_{r}(J[\mathfrak{p}])/\{\pm 1\}$ consisting of automorphisms of $J[\mathfrak{p}]$ which are symplectic with respect to $(e_{J,7})^{r}$. Therefore, the field of definition of the automorphisms of $X_{J[\mathfrak{p}]}^{\pm}(7)$ is equal to $\mathbb{Q}(\mathcal{K}[\mathfrak{p}])$ and $X_{J[\mathfrak{p}]}^{\pm}(7)$ admits no non-trivial automorphisms over $L_{1}L_{2}$ that are not defined over $\mathbb{Q}$. Suppose there is such an automorphism $\tau$ defined over $\mathbb{Q}$. Since the mod $7$ Galois representation attached to $J[\mathfrak{p}]$ is surjective ($J[\mathfrak{p}]$ is isomorphic over $L_{i}$ to a quadratic twist of $E_{i}[7]$ for each $i=1,2$) the element $\tau$ is contained in the centre of $\operatorname{Aut}_{r}(J[\mathfrak{p}])/\{\pm 1\}$ which is isomorphic to $\operatorname{PSL}_{2}(\mathbb{Z}/7\mathbb{Z})$ as an abstract group. Therefore $\tau$ is the identity and the claim follows from Lemma 4.3. ∎
5. Proving $(7,\mathfrak{p})$-congruences and Theorem 1.1
We now prove Theorem 1.1. In order to apply visibility we must first show that the pairs $(C,E)$ in Table 1 are in fact $(7,\mathfrak{p})$-congruent (not simply up to quadratic twist, as we proved in Section 4).
Lemma 5.1.
Let $E/\mathbb{Q}$ be an elliptic curve and let $J/\mathbb{Q}$ be a genus $2$ Jacobian with RM by $\mathcal{O}_{D}$. Suppose that $(p)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathcal{O}_{D}$ and that there exists a squarefree integer $d\in\mathbb{Z}$ such that $E^{d}$ and $J$ are $(p,\mathfrak{p})$-congruent. Then $d$ is supported on the set of primes consisting of $p$, the bad primes of $E$, and the bad prime of $J$.
Proof.
This is similar to [FK22, Proposition 4.18] and [Fre24, Lemma 3.6] (see also [Ser72, Lemme 8]). Let $\ell\neq p$ be a prime at which $J$ has good reduction and at which $E$ has potentially good reduction. Let $\mathbb{Q}_{\ell}^{\textrm{ur}}$ be the maximal unramified extension of $\mathbb{Q}_{\ell}$ and let $K=\mathbb{Q}_{\ell}^{\mathrm{ur}}(J[p])$. By [ST68, §2 Corollary 3] if $A/\mathbb{Q}_{\ell}$ is an abelian variety with potential good reduction at $\ell$, then for each $p\neq\ell$ the field $\mathbb{Q}_{\ell}^{\mathrm{ur}}(A[p])$ is the smallest extension of $\mathbb{Q}_{\ell}^{\mathrm{ur}}$ over which $A$ attains good reduction. But then we have $\mathbb{Q}_{\ell}^{\mathrm{ur}}(E[p])=\mathbb{Q}_{\ell}^{\mathrm{ur}}(J[\mathfrak{p}])\subset K=\mathbb{Q}_{\ell}^{\mathrm{ur}}$, as required. ∎
Proposition 5.2.
For each pair $(E,C)$ of elliptic curve $E/\mathbb{Q}$ and genus 2 curve $C/\mathbb{Q}$ in Table 1 we have a $(7,\mathfrak{p})$-congruence between $E$ and $J=\operatorname{Jac}(C)$ for some choice of $\mathfrak{p}|7$ in $\mathcal{O}_{D}$.
Proof.
Let $\ell\neq 7$ be a good prime for $C$ and $E$. By [Fis16, (5.2)] (which follows from [FLS+01, Section 2.1] or [MS93, Lemma 3]) a $(7,\mathfrak{p})$-congruence between $E^{d}/\mathbb{Q}$ and $J/\mathbb{Q}$ gives a congruence modulo $7$
(2) $a_{\ell}(E^{d})^{2}-t_{\ell}a_{\ell}(E^{d})+n_{\ell}\equiv 0\pmod{7}$
where $t_{\ell}=\ell+1-N_{1}$ and $n_{\ell}=(N_{1}^{2}+N_{2})/2-(\ell+1)N_{1}-\ell$ where $N_{1}=\#C(\mathbb{F}_{\ell})$ and $N_{2}=\#C(\mathbb{F}_{\ell^{2}})$.
Testing (2) on the divisors $d$ of the product $7$ and the bad primes of $E$ and $C$ shows that $E^{d}$ and $J$ are not $(7,\mathfrak{p})$-congruent for any $d\neq 1$ (by Lemma 5.1). By 4.4 $E$ and $J$ are $(7,\mathfrak{p})$-congruent up to quadratic twist since $E$ corresponds to a point on one of the twists $X_{J[\mathfrak{p}]}^{\pm}(7)$. It therefore follows that $E$ and $J$ are $(7,\mathfrak{p})$-congruent for some choice of $\mathfrak{p}$ dividing $7$. ∎
Using the congruences supplied by 5.2 we now prove Theorems 1.1 and 1.5 by applying [Fis16, Theorem 2.2].
Proof of Theorems 1.1 and 1.5.
This follows from [Fis16, Theorem 2.2], as we detail below.
Let $E/\mathbb{Q}$ and $J/\mathbb{Q}$ be one of the pairs of elliptic curve and genus $2$ Jacobian from Theorem 1.1 or 1.5. We check that in each case $J$ is geometrically simple by applying the condition in [CF96, Section 14.4] and [Sto95]. The $7$-torsion subgroups of $E(\mathbb{Q})$ and $J(\mathbb{Q})$ are trivial. The rank of $E/\mathbb{Q}$ is $2$ and the rank of $J/\mathbb{Q}$ is 00 (the rank of $J/\mathbb{Q}$ is bounded using $2$-descent, which is implemented as RankBounds in Magma). For each discriminant $D$ appearing in Theorems 1.1 and 1.5 the prime $7$ not only splits in $\mathcal{O}_{D}$, but $7=\operatorname{Nm}\eta$ for some $\eta\in\mathcal{O}_{D}$. In particular the isogeny $\mathfrak{p}$ is equal to the multiplication-by-$\eta$-map on $J$, and $J(\mathbb{Q})/\mathfrak{p}J(\mathbb{Q})=0$.
By 5.2 the elliptic curve $E$ is $(7,\mathfrak{p})$-congruent to $J$. The abelian varieties $E$ and $J$ have good reduction at $7$, so by [Fis16, Theorem 2.2] it suffices to show that the Tamagawa numbers of $E/\mathbb{Q}$ and $J/\mathbb{Q}$ are coprime to $7$.
We compute the Tamagawa numbers of $E/\mathbb{Q}$ using Magma. Except for the Jacobian of conductor $3200^{2}$ in Table 1, for each bad prime $p$ of $J$ one may check that the order of the geometric component group of $J$ at $p$ is coprime to $7$ using Liu’s genus2reduction in SageMath and Donnelly’s Magma functions RegularModel and ComponentGroup.
For the Jacobian of conductor $3200^{2}$ in Table 1 the computation of the Tamagawa number of $J/\mathbb{Q}$ at $2$ was carried out in the appendix to [KS23] (where it is shown that the Tamagawa number is $1$). ∎
6. Evidence towards 1.6
Consider the genus $2$ curve $C/\mathbb{Q}$ with LMFDB label 385641.a.385641.1 and Weierstrass equation
$C:y^{2}+(x^{3}+1)y=-6x^{4}+6x^{3}+27x^{2}-30x-22.$
The Jacobian $J/\mathbb{Q}$ of $C$ has RM by $\mathbb{Z}[\sqrt{2}]$. In [KS23, A.3] it is noted that the Birch and Swinnerton-Dyer conjecture predicts $|\Sha(J^{-11}/\mathbb{Q})|=7^{2}$. By 4.4, for some choice of factorisation $(7)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathbb{Z}[\sqrt{2}]$ we have models
$X_{J[\mathfrak{p}]}^{\pm}(7):$ $\;-2x_{0}^{4}+39x_{0}^{3}x_{1}+11x_{0}^{3}x_{2}-42x_{0}^{2}x_{1}^{2}-18x_{0}^{2}x_{1}x_{2}+20x_{0}x_{1}^{3}-6x_{0}x_{1}^{2}x_{2}$
$+12x_{0}x_{1}x_{2}^{2}-7x_{0}x_{2}^{3}-24x_{1}^{4}+13x_{1}^{3}x_{2}+15x_{1}^{2}x_{2}^{2}+9x_{1}x_{2}^{3}+x_{2}^{4}=0,$
$X_{J[\mathfrak{p}]}^{\mp}(7):$  $\;2x_{0}^{4}+5x_{0}^{3}x_{1}+9x_{0}^{3}x_{2}+6x_{0}^{2}x_{2}^{2}-x_{0}x_{1}^{3}-6x_{0}x_{1}^{2}x_{2}+12x_{0}x_{1}x_{2}^{2}+2x_{0}x_{2}^{3}$
$-x_{1}^{4}-3x_{1}^{3}x_{2}+3x_{1}^{2}x_{2}^{2}+17x_{1}x_{2}^{3}+12x_{2}^{4}=0,$
$X_{J[\bar{\mathfrak{p}}]}^{\pm}(7):$ $\;x_{0}^{4}-3x_{0}^{3}x_{1}-28x_{0}^{3}x_{2}-15x_{0}^{2}x_{1}^{2}-3x_{0}^{2}x_{1}x_{2}+39x_{0}^{2}x_{2}^{2}-6x_{0}x_{1}^{3}-12x_{0}x_{1}^{2}x_{2}$
$-6x_{0}x_{1}x_{2}^{2}-29x_{0}x_{2}^{3}+3x_{1}^{4}+9x_{1}^{3}x_{2}+30x_{1}^{2}x_{2}^{2}-3x_{1}x_{2}^{3}-10x_{2}^{4}=0,$
$X_{J[\bar{\mathfrak{p}}]}^{\mp}(7):$ $\;-4x_{0}^{4}+6x_{0}^{3}x_{1}+7x_{0}^{3}x_{2}+3x_{0}^{2}x_{1}^{2}+12x_{0}x_{1}^{3}+6x_{0}x_{1}^{2}x_{2}-9x_{0}x_{1}x_{2}^{2}-x_{0}x_{2}^{3}$
$-6x_{1}^{4}-3x_{1}^{3}x_{2}+3x_{1}^{2}x_{2}^{2}+6x_{1}x_{2}^{3}+x_{2}^{4}=0.$
We were unable to find rational points on any of these curves, except on $X_{J[\mathfrak{p}]}^{\mp}(7)$ where we find exactly one point which corresponds to the elliptic curve $E/\mathbb{Q}$ with LMFDB label 1242.m1 and Weierstrass equation $y^{2}+xy+y=x^{3}-x^{2}-1666739x-2448131309$. Using the argument in 5.2 it can be shown that $E$ and $J$ are $(7,\mathfrak{p})$-congruent. However, the quadratic twist of $E$ by $-11$ has trivial Mordell-Weil group, so cannot be used to visualise the (conjectural) non-trivial elements of $\Sha(J^{-11}/\mathbb{Q})[7]$.

### v2

@@
 Abstract.
-We describe an algorithm which, on input a genus $2$ curve $C/\mathbb{Q}$ whose Jacobian $J/\mathbb{Q}$ has real multiplication by a quadratic order in which $7$ splits, outputs twists of the Klein quartic curve parametrising elliptic curves whose mod $7$ Galois representations are isomorphic to a sub-representation of the mod $7$ Galois representation attached to $J/\mathbb{Q}$.
+Let $C/\mathbb{Q}$ be a genus $2$ curve whose Jacobian $J/\mathbb{Q}$ has real multiplication by a quadratic order in which $7$ splits.
+We describe an algorithm which outputs twists of the Klein quartic curve which parametrise elliptic curves whose mod $7$ Galois representations are isomorphic to a sub-representation of the mod $7$ Galois representation attached to $J/\mathbb{Q}$.
 Applying this algorithm to genus $2$ curves of small conductor in families of Bending and Elkies–Kumar we exhibit a number of genus $2$ Jacobians whose Tate–Shafarevich groups (unconditionally) contain a non-trivial element of order $7$ which is visible in an abelian three-fold.
@@
 In this article we prove the following theorem.
-Theorem 1.1.
+Theorem 1.
 […]
-Remark 1.2.
-The genus $2$ Jacobian $J/\mathbb{Q}$ of conductor $3200^{2}$ in Table 1 was included in an appendix to [KS23] joint with Keller and Stoll where the strong Birch and Swinnerton-Dyer conjecture is also verified for $J$.
+Remark 2.
+The genus $2$ Jacobian $J/\mathbb{Q}$ of conductor $3200^{2}$ in Table 1.1 was included in an appendix to [KS25] joint with Keller and Stoll where the strong Birch and Swinnerton-Dyer conjecture is also verified for $J$.
 $D$ $f(x)$ $\sqrt{N_{J}}$ $E$
-$8$ $-10(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 3200 3200.a1
-$8$ $165(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 39325 39325.c1
-$37$ $-13(27x^{6}-54x^{5}-90x^{4}+228x^{3}+15x^{2}-90x-23)$ 73008 73008.n1
-$8$ $-51(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 93925 93925.d1
-$8$ $285(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 117325 117325.c1
-$8$ $-46(x^{6}+6x^{5}-20x^{4}+240x^{3}+70x^{2}-84x+12)$ 203136 203136.i2
-$8$ $-5(3x^{6}+12x^{5}+89x^{4}-56x^{3}-7x^{2}-132x+99)$ 211200 211200.c1
-$8$ $465(9x^{6}-12x^{5}+64x^{4}-56x^{3}+136x^{2}-60x+84)$ 216225 432450.ci1
-$8$ $-30(11x^{6}-18x^{5}+47x^{4}+6x^{3}+71x^{2}+18x+27)$ 244800 244800.dc1
-$8$ $-390(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 270400 270400.dc2
-$8$ $-177(3x^{6}+12x^{4}-10x^{3}-12x+11)$ 281961 2819610*
-$8$ $-22(x^{6}-24x^{5}+100x^{4}+102x^{3}-80x^{2}-132x-39)$ 302016 302016.p1
-$8$ $-6(13x^{6}-116x^{5}-316x^{4}+58x^{3}+264x^{2}-116x+13)$ 313920 313920.bb1
-$8$ $-55(13x^{6}-116x^{5}-316x^{4}+58x^{3}+264x^{2}-116x+13)$ 329725 659450*
-$8$ $-110(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 387200 8905600*
-$8$ $11(x^{6}+6x^{5}+11x^{4}-13x^{2}+6x-2)$ 423984 423984.by1
+88 $10(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 3200 3200.a1
+88 $165(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 39325 39325.c1
+$37$ $13(27x^{6}-54x^{5}-90x^{4}+228x^{3}+15x^{2}-90x-23)$ 73008 73008.n1
+88 $51(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 93925 93925.d1
+88 $285(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 117325 117325.c1
+88 $62(9x^{6}-12x^{5}+64x^{4}-56x^{3}+136x^{2}-60x+84)$ 184512 184512.bw1
+88 $46(x^{6}+6x^{5}-20x^{4}+240x^{3}+70x^{2}-84x+12)$ 203136 203136.i2
+88 $5(3x^{6}+12x^{5}+89x^{4}-56x^{3}-7x^{2}-132x+99)$ 211200 211200.c1
+88 $465(9x^{6}-12x^{5}+64x^{4}-56x^{3}+136x^{2}-60x+84)$ 216225 432450.ci1
+88 $30(11x^{6}-18x^{5}+47x^{4}+6x^{3}+71x^{2}+18x+27)$ 244800 244800.dc1
+88 $13(x^{6}-2x^{5}+3x^{4}x^{3}-7x^{2}-2x+1)$ 256880 51376.e1
+88 $390(x^{6}+6x^{5}+27x^{4}-2x^{3}+45x^{2}+20)$ 270400 270400.dc2
+88 $177(3x^{6}+12x^{4}-10x^{3}-12x+11)$ 281961 2819610*
+88 $22(x^{6}-24x^{5}+100x^{4}+102x^{3}-80x^{2}-132x-39)$ 302016 302016.p1
+88 $6(13x^{6}-116x^{5}-316x^{4}+58x^{3}+264x^{2}-116x+13)$ 313920 313920.bb1
+88 $55(13x^{6}-116x^{5}-316x^{4}+58x^{3}+264x^{2}-116x+13)$ 329725 659450*
+88 $110(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 387200 8905600*
+88 $11(x^{6}+6x^{5}+11x^{4}-13x^{2}+6x-2)$ 423984 423984.by1
+Table $1.$ Examples of genus $2$ curves $C\colon y^{2}=f(x)$ whose Jacobians $J/\mathbb{Q}$ have conductor $N_{J}<(500\,000)^{2}$, and such that $\Sha(J/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$ (we do not claim, nor expect, this list to be complete).
+The Jacobians have real multiplication by the quadratic order $\mathcal{O}_{D}$ of discriminant $D$ and the subgroup of $\Sha(J/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence between $E/\mathbb{Q}$ and $J/\mathbb{Q}$ where $\mathfrak{p}$ divides 77 in $\mathcal{O}_{D}$.
+We write NN* for an elliptic curve of conductor NN which does not appear in the LMFDB (i.e., if N>500  [^0] N>500\,000).
+Explicit Weierstrass equations for the corresponding elliptic curves are given in [Fre].
+The curves $C/\mathbb{Q}$ were generated from [Ben99, Theorem 4.1] and [EK14, CFM24].
+Conductors were computed using [DD19].
 […]
-The group $\Sha(A/K)$ is a torsion group and is conjectured to be finite.
-It is conjectured that for every prime number $p>0$ and each integer $g>0$ there exists an absolutely simple abelian variety (i.e., one which is not isogenous over $\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}\mkern-1.5mu}\mkern 1.5mu$ to a product) of dimension $g$ for which $\Sha(A/\mathbb{Q})[p]\neq 0$.
-Indeed, it is even conjectured that $\Sha(A^{d}/\mathbb{Q})[p]\neq 0$ for a positive proportion of quadratic twists of a fixed abelian variety $A/\mathbb{Q}$ (see e.g., [BKLS21, Conjecture 1.1]).
 In spite of this, for general values of $g$ and $p$, constructing an example of an absolutely simple $g$-dimensional abelian variety $A/\mathbb{Q}$ with an $p$-torsion element contained in $\Sha(A/\mathbb{Q})$ is an open problem.
 By allowing the dimension of $A$ to increase with $p$, Shnidman and Weiss [SW22] construct examples of absolutely simple abelian varieties with $\Sha(A/\mathbb{Q})[p]\neq 0$.
-Flynn and Shnidman [FS22] extended this result to show $\Sha(A/\mathbb{Q})[p]$ can be arbitrarily large.
+Flynn and Shnidman [FS25] extended this result to show $\Sha(A/\mathbb{Q})[p]$ can be arbitrarily large.
 When $A$ has dimension $2$, Bruin, Flynn, and Testa [BFT14, Fly15] found examples of absolutely simple genus $2$ Jacobians with $3$ and $5$-torsion in their Tate–Shafarevich groups.
@@
 Let $K$ be a number field and let $A/K$ and $A^{\prime}/K$ be abelian varieties equipped with isogenies $\psi\colon A\to B$ and $\psi^{\prime}\colon A^{\prime}\to B^{\prime}$.
-Definition 1.3.
+Definition 3.
 We say that $A/K$ and $A^{\prime}/K$ are $(\psi,\psi^{\prime})$-congruent if there exists a $G_{K}$-equivariant group isomorphism $\phi\colon A[\psi]\to A^{\prime}[\psi^{\prime}]$.
@@
 More precisely, if $B(K)/\psi A(K)=0$ then, under mild hypotheses applied at the bad primes of $A$ and $A^{\prime}$ and the primes dividing $|\Delta|$ (see [Fis16, Theorem 2.2]), the group $\operatorname{Vis}_{Z}\Sha(A/K)$ of elements of $\Sha(A/K)$ that are visible in $Z=(A\times A^{\prime})/\Delta$ is equal to $B^{\prime}(K)/\psi^{\prime}A^{\prime}(K)$.
-The central idea for proving Theorem 1.1 is to construct examples of genus $2$ curves $C/\mathbb{Q}$ with the property that there exists a $(7,\psi)$-congruence between an elliptic curve $E/\mathbb{Q}$ and the Jacobian $J=\operatorname{Jac}(C)$ of $C$, for some isogeny $\psi\colon J\to B$.
+The central idea for proving Theorem 1 is to construct examples of genus $2$ curves $C/\mathbb{Q}$ with the property that there exists a $(7,\psi)$-congruence between an elliptic curve $E/\mathbb{Q}$ and the Jacobian $J=\operatorname{Jac}(C)$ of $C$, for some isogeny $\psi\colon J\to B$.
 Assuming that the local conditions are satisfied, it then suffices to show that $B(\mathbb{Q})/\psi J(\mathbb{Q})=0$ and that $E(\mathbb{Q})/7E(\mathbb{Q})\neq 0$ (which in practice is computationally less intensive than performing a $\psi$-descent on $J$).
@@
 Writing $(7)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathcal{O}_{D}$ by abuse of notation we write $\mathfrak{p}\colon J\to B$ for the isogeny with kernel consisting of those $P\in J(\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}\mkern-1.5mu}\mkern 1.5mu)$ annihilated by $\mathfrak{p}$.
-In this case, $\ker\mathfrak{p}$ is isomorphic as a group to $(\mathbb{Z}/7\mathbb{Z})^{2}$ and under suitable hypotheses (see Lemma 2.2) comes equipped with a natural alternating bilinear pairing.
+In this case, $\ker\mathfrak{p}$ is isomorphic as a group to $(\mathbb{Z}/7\mathbb{Z})^{2}$ and under suitable hypotheses (see Lemma 8) comes equipped with a natural alternating bilinear pairing.
 In Section 3.2 we describe an algorithm for determining (a $q$-adic approximation to) a pair of twists of the Klein quartic (the modular curve $X(7)$) which parametrise elliptic curves that are $(7,\mathfrak{p})$-congruent to a fixed genus $2$ Jacobian $J/\mathbb{Q}$ with real multiplication by $\mathcal{O}_{D}$ (our algorithm is subject to the technical hypothesis that $J[\mathfrak{p}]$ is an irreducible $G_{\mathbb{Q}}$-module).
-Remark 1.4.
-Since abelian surfaces $J/\mathbb{Q}$ with RM by $\mathcal{O}_{D}$ are modular (this follows from Serre’s conjecture [KW09a, KW09b]) we may associate to $J$ a weight $2$ newform with coefficients in $\mathcal{O}_{D}$ and level $\sqrt{N_{J}}$, where $N_{J}$ is the conductor of $J$ (in particular $N_{J}$ is a perfect square).
+Remark 4.
+Since abelian surfaces $J/\mathbb{Q}$ with RM by $\mathcal{O}_{D}$ are modular (this follows from Serre’s conjecture [KW09a, KW09b]) we may associate to $J$ a weight 22 newform with coefficients in $\mathcal{O}_{D}$ and level $\sqrt{N_{J}}$, where $N_{J}$ is the conductor of $J$ (in particular $N_{J}$ is a perfect square).
 We compute these twists of $X(7)$ for examples of genus $2$ Jacobians of small conductor provided by the real multiplication families of Bending [Ben98, Ben99] and of Elkies–Kumar [EK14, CFM24].
 By searching for rational points on these twists, we find a number of putative examples of $(7,\mathfrak{p})$-congruences between an elliptic curve $E/\mathbb{Q}$ and a genus $2$ Jacobian $J/\mathbb{Q}$.
-Adapting an approach of Fisher [Fis16, Section 6] we prove these congruences in 5.2.
+Adapting an approach of Fisher [Fis16, Section 6] we prove these congruences in 22.
 For an abelian variety $A/\mathbb{Q}$ we write $A^{d}/\mathbb{Q}$ for the quadratic twist of $A$ by a squarefree integer $d\in\mathbb{Z}$.
@@
 [FK22, Lemma 4.15]).
-To construct the examples in Theorem 1.1 we simply search for quadratic twists of the examples in 5.2 where there is a rank discrepancy between $E^{d}/\mathbb{Q}$ and $J^{d}/\mathbb{Q}$.
-In addition to Theorem 1.1 we prove that there exist examples of such genus $2$ Jacobians with $7$-torsion in their Tate–Shafarevich groups and with real multiplication by $\mathcal{O}_{D}$ for several fundamental discriminants $D>0$.
-Theorem 1.5.
+To construct the examples in Theorem 1 we simply search for quadratic twists of the examples in 22 where there is a rank discrepancy between $E^{d}/\mathbb{Q}$ and $J^{d}/\mathbb{Q}$.
+In addition to Theorem 1 we prove that there exist examples of such genus $2$ Jacobians with $7$-torsion in their Tate–Shafarevich groups and with real multiplication by $\mathcal{O}_{D}$ for several fundamental discriminants $D>0$.
+Theorem 5.
 For each $D=8$, $29$, $37$, $44$, and $57$ there exists an absolutely simple genus $2$ Jacobian $J/\mathbb{Q}$ with real multiplication by $\mathcal{O}_{D}$ such that $\Sha(J/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$.
-Examples are furnished by the Jacobians of the curves $C\colon y^{2}=f(x)$ given in Table 2.
+Examples are furnished by the Jacobians of the curves $C\colon y^{2}=f(x)$ given in Table 1.2.
 $D$ $f(x)$ $\sqrt{N_{J}}$ $E$
-$8$ $-10(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 3200 3200.a1
-$29$ $-2470(8x^{6}-2x^{5}+68x^{4}+221x^{3}+122x^{2}+986x+1588)$ 40019200 760364800*
-$37$ $-39(x^{6}-45x^{4}-68x^{3}+504x^{2}+180x-1193)$ 73008 73008.n1
-$44$ $-39(14x^{6}-30x^{5}+85x^{4}+700x^{3}-1325x^{2}+3000x+18000)$ 608400 608400*
+88 $10(x^{6}-4x^{5}-3x^{4}+8x^{3}+25x^{2}+20x+5)$ 3200 3200.a1
+$29$ $2470(8x^{6}-2x^{5}+68x^{4}+221x^{3}+122x^{2}+986x+1588)$ 40019200 760364800*
+$37$ $39(x^{6}-45x^{4}-68x^{3}+504x^{2}+180x-1193)$ 73008 73008.n1
+$44$ $39(14x^{6}-30x^{5}+85x^{4}+700x^{3}-1325x^{2}+3000x+18000)$ 608400 608400*
 $57$ $1479(80x^{6}+279x^{4}+186x^{3}+243x^{2}+324x+108)$ 590609070 7677917910*
-Table 2.
+Table 1.2.
 Examples of genus $2$ curves $C\colon y^{2}=f(x)$ whose Jacobians $J/\mathbb{Q}$ have real multiplication by the quadratic order $\mathcal{O}_{D}$ of discriminant $D$ and such that $\Sha(J/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$.
-The subgroup of $\Sha(J/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence between $E/\mathbb{Q}$ and $J/\mathbb{Q}$ where $\mathfrak{p}$ divides $7$ in $\mathcal{O}_{D}$.
+The subgroup of $\Sha(J/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence between $E/\mathbb{Q}$ and $J/\mathbb{Q}$ where $\mathfrak{p}$ divides 77 in $\mathcal{O}_{D}$.
+We write NN* for an elliptic curve of conductor NN which does not appear in the LMFDB (i.e., if N>500  [^0] N>500\,000).
 Explicit Weierstrass equations for the corresponding elliptic curves are given in [Fre].
@@
 Conductors were computed using [DD19].
-In [KS23, A.3] it is observed that the Birch and Swinnerton–Dyer conjecture predicts that $\Sha(J^{-11}/\mathbb{Q})$ contains a subgroup isomoprhic to $(\mathbb{Z}/7\mathbb{Z})^{2}$, where $J/\mathbb{Q}$ is the Jacobian of the genus $2$ curve with LMFDB label 385641.a.385641.1.
+In [KS25, A.3] it is observed that the Birch and Swinnerton–Dyer conjecture predicts that $\Sha(J^{-11}/\mathbb{Q})$ contains a subgroup isomorphic to $(\mathbb{Z}/7\mathbb{Z})^{2}$, where $J/\mathbb{Q}$ is the Jacobian of the genus $2$ curve with LMFDB label 385641.a.385641.1.
 Since $J$ has RM by $\mathcal{O}_{8}$ it is natural to ask whether the $7$-torsion in $\Sha(J^{-11}/\mathbb{Q})$ is made visible by a $(7,\mathfrak{p})$-congruence with an elliptic curve.
 By computing the relevant twists of $X(7)$ we give evidence that this is not the case (see Section 6).
-Conjecture 1.6.
+Conjecture 6.
 Let $C/\mathbb{Q}$ be the genus $2$ curve with LMFDB label 385641.a.385641.1 and let $J/\mathbb{Q}$ be its Jacobian.
@@
 It takes as input a genus $2$ Jacobian $J/\mathbb{Q}$ with RM by an order in which $7$ splits, and outputs four twists of $X(7)$ which parametrise elliptic curves $(7,\mathfrak{p})$-congruent to $J/\mathbb{Q}$.
-The outputs of the algorithm in Section 3.2 are not guaranteed to be correct, however in Section 4 we prove that the output is correct in many cases (for example for the curves in Tables 1 and 2).
+The outputs of the algorithm in Section 3.2 are not guaranteed to be correct, however in Section 4 we prove that the output is correct in many cases (for example for the curves in Tables 1.1 and 1.2).
 In particular, in Section 4 we prove that the twists we obtain are isomorphic to those which parametrise elliptic curves $(7,\mathfrak{p})$-congruent to $J$.
-In Section 5 we prove Theorems 1.1 and 1.5 by proving that the pairs $(E,J)$ in Tables 1 and 2 are $(7,\mathfrak{p})$-congruent, and by checking that the local hypotheses in [Fis16, Theorem 2.2] are satisfied.
+In Section 5 we prove Theorems 1 and 5 by proving that the pairs $(E,J)$ in Tables 1.1 and 1.2 are $(7,\mathfrak{p})$-congruent, and by checking that the local hypotheses in [Fis16, Theorem 2.2] are satisfied.
 Finally in Section 6 we give explicit examples of the Klein quartic twists for the Jacobian of the genus $2$ curve with LMFDB label 385641.a.385641.1.
-By searching for rational points on these twists, we give evidence towards 1.6.
+By searching for rational points on these twists, we give evidence towards 6.
 1.2. Acknowledgements
@@
 In the case when $p=7$ the curve $X(7)$ is the Klein quartic [Kle78] (see [Elk99] for a detailed discussion).
-[…]
+Let $K$ be a field of characteristic zero.
+A symplectic abelian group over $K$ is a pair $(M,e_{M})$ where $M$ is a $G_{K}$-module equipped with a ($G_{K}$-equivariant) alternating, bilinear pairing $e_{M}\colon M\times M\to\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu^{\times}$.
+We equip $\mu_{p}\times\mathbb{Z}/p\mathbb{Z}$ with the natural alternating pairing $\langle(\zeta,n),(\xi,m)\rangle=\zeta^{m}\xi^{-n}$.
 2.1. The modular curve $X(p)$
@@
 As an abstract group $\Gamma\!_{p}$ is isomorphic to $\operatorname{SL}_{2}(\mathbb{Z}/p\mathbb{Z})$, but it comes equipped with a non-trivial action of $\operatorname{Gal}(\mathbb{Q}(\mu_{p})/\mathbb{Q})$.
-The matrix ±$\pm I$ acts trivially on $Y(p)$ and therefore the action of $\Gamma\!_{p}$ factors through $\Gamma\!_{p}/\{\pm I\}$.
+The matrix $\pm I$ acts trivially on $Y(p)$ and therefore the action of $\Gamma\!_{p}$ factors through $\Gamma\!_{p}/\{\pm I\}$.
 This action extends to an action on $X(p)$, and the quotient realises the forgetful morphism $X(p)\to X(1)$ given by taking $j$-invariants.
@@
 The following lemma is well known and follows by construction (cf. [PSS07, Section 4.4]).
-Lemma 2.1.
+Lemma 7.
 For each field $L/K$ the $L$-rational points on $X_{M}^{r}(p)$ correspond to pairs $(E,\phi)$ where $E/L$ is an elliptic curve and $\phi\colon M\cong E[p]$ is an isomorphism of $G_{L}$-modules for which $e_{E,p}(\phi(P),\phi(Q))=e_{M}(P,Q)^{r}$ for each $P,Q\in M$.
@@
 The choice of principal polarisation $\lambda$ induces the alternating, bilinear $p$-Weil pairing $e_{J,p}\colon J[p]\times J[p]\to\mu_{p}$.
-Lemma 2.2.
+Lemma 8.
 Let $J/K$ be a genus $2$ Jacobian with RM by $\mathcal{O}_{D}$.
@@
 This is [CFM24, Lemma 3.4], cf. the proof of [Fis16, Proposition 6.1] when $\mathcal{O}_{D}=\mathbb{Z}[\sqrt{2}]$. ∎
-Note that by Lemma 2.2 we may define the twists $X_{J[\mathfrak{p}]}^{\pm}(p)$ and $X_{J[\bar{\mathfrak{p}}]}^{\pm}(p)$ which (by Lemma 2.1) parametrise elliptic curves which are $(p,\mathfrak{p})$-congruent (respectively $(p,\bar{\mathfrak{p}})$-congruent) to the genus $2$ Jacobian $J/K$.
+Note that by Lemma 8 we may define the twists $X_{J[\mathfrak{p}]}^{\pm}(p)$ and $X_{J[\bar{\mathfrak{p}}]}^{\pm}(p)$ which (by Lemma 7) parametrise elliptic curves which are $(p,\mathfrak{p})$-congruent (respectively $(p,\bar{\mathfrak{p}})$-congruent) to the genus $2$ Jacobian $J/K$.
 2.4. Explicit twisting for $X(7)$
-[…]
+Following [Fis14] we give an explicit description for the twists $X_{M}^{\pm}(7)$ as plane quartic curves.
+Recall that we write $\Gamma\!_{7}$ for the automorphism group (scheme) of the symplectic abelian group $\mu_{7}\times\mathbb{Z}/7\mathbb{Z}$.
+Following Klein [Kle78, Elk99, Fis14] consider the representation $\operatorname{SL}_{2}(\mathbb{Z}/7\mathbb{Z})\to\operatorname{GL}_{3}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu)$ which maps the generators $S=\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right)$ and $T=\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right)$ to
 $\frac{1}{\sqrt{-7}}\begin{pmatrix}\zeta_{7}-\zeta_{7}^{6}&\zeta_{7}^{2}-\zeta_{7}^{5}&\zeta_{7}^{4}-\zeta_{7}^{3}\\ \zeta_{7}^{2}-\zeta_{7}^{5}&\zeta_{7}^{4}-\zeta_{7}^{3}&\zeta_{7}-\zeta_{7}^{6}\\ \zeta_{7}^{4}-\zeta_{7}^{3}&\zeta_{7}-\zeta_{7}^{6}&\zeta_{7}^{2}-\zeta_{7}^{5}\\ \end{pmatrix}\quad\text{and}\quad\begin{pmatrix}\zeta_{7}&0&0\\ 0&\zeta_{7}^{4}&0\\ 0&0&\zeta_{7}^{2}\\ \end{pmatrix}$.
-Composing with the natural $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-isomorphism $\Gamma\!_{7}\cong\operatorname{SL}_{2}(\mathbb{Z}/7\mathbb{Z})$ gives a $G_{K}$-equivariant homomorphism $\rho\colon\Gamma\!_{7}\to\operatorname{GL}_{2}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu)$.
+Composing with the natural $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-isomorphism $\Gamma\!_{7}\cong\operatorname{SL}_{2}(\mathbb{Z}/7\mathbb{Z})$ gives a $G_{K}$-equivariant homomorphism $\rho\colon\Gamma\!_{7}\to\operatorname{GL}_{3}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu)$.
 When $\Gamma\!_{7}$ acts on $\mathbb{P}^{2}$ via $\rho$} the Klein quartic curve $X(7)\subset\mathbb{P}^{2}$ given by
@@
 By Hilbert’s Theorem 90 the cocycle $\xi$ is a coboundary, that is, there exists a matrix $A\in\operatorname{GL}_{3}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu)$ such that $\sigma(A^{-1})A=\xi(\sigma)$ for every $\sigma\in G_{K}$.
-Lemma 2.3.
+Lemma 9.
 The curves $X_{M}(7)$ and $X_{M}^{-}(7)$ are isomorphic to the images of the morphisms $X(7)\to\mathbb{P}^{2}$ given by
@@
 Finally, we note that it is simple to recover the moduli interpretation for a given twist of $X(7)$ following [Fis14, Section 4.1] (cf. [PSS07, Section 7.1]).
-Lemma 2.4.
+Lemma 10.
 Let $\mathcal{X}/K$ be a twist of the Klein quartic given by the vanishing of a homogeneous quartic polynomial $\mathcal{F}(x_{0},x_{1},x_{2})$ in $\mathbb{P}^{2}$.
@@
 3.1. Hilbert’s Theorem 90 is effective
-To compute twists of $X(7)$ using Lemma 2.3 we will need to compute matrices which realise a given cocycle as a coboundary.
+To compute twists of $X(7)$ using Lemma 9 we will need to compute matrices which realise a given cocycle as a coboundary.
 Towards this, note that the standard proof of Hilbert’s Theorem 90 is “nearly” effective.
@@
 The proof therefore suggests an algorithm.
-Algorithm 3.1 Generating matrices which realise Hilbert’s Theorem 90.
+Algorithm 11 Generating matrices which realise Hilbert’s Theorem 90.
 Input: A cocycle $\xi\in H^{1}(\operatorname{Gal}(L/K),\operatorname{GL}_{n}(L))$.
@@
 7:end if
-Remark 3.2.
-In step 1 of Algorithm 3.1 the user must choose a matrix $c\in\operatorname{GL}_{n}(L)$.
+Remark 12.
+In step 1 of Algorithm 11 the user must choose a matrix $c\in\operatorname{GL}_{n}(L)$.
 In our application the extension $L/\mathbb{Q}$ will be a finite extension and we will have access to a LLL-reduced $\mathbb{Z}$-basis $\{a_{1},...,a_{m}\}$ for the ring of integers $\mathcal{O}_{L}\subset L$.
@@
 This approach extends to choosing a matrix $c\in\operatorname{GL}_{n}(L)$ by choosing the $n^{2}$ entries as described (in practice, we also choose almost all $x_{i}$ to be equal to zero).
-Remark 3.3.
-In principle the recursive step in Algorithm 3.1 may be called many times. In practice, however, we have found very few instances when Algorithm 3.1 fails to terminate in one iteration.
+Remark 13.
+In principle the iteration in Step 6 in Algorithm 11 may be called many times. In practice, however, we have found very few instances when Algorithm 11 fails to terminate in one iteration.
 3.2. The main algorithm
@@
   5. 5:
-    We compute (a $\mathfrak{q}$-adic approximation to) a matrix $A\in\operatorname{GL}_{3}(L)\subset\operatorname{GL}_{3}(L_{\mathfrak{q}})$ which realises $\xi$ as a coboundary, using Algorithm 3.1.
+    We compute (a $\mathfrak{q}$-adic approximation to) a matrix $A\in\operatorname{GL}_{3}(L)\subset\operatorname{GL}_{3}(L_{\mathfrak{q}})$ which realises $\xi$ as a coboundary, using Algorithm 11.
   6. 6:
     We twist $X(7)$ by $A$ and $A^{-T}$ to obtain curves $\mathcal{X}^{\pm}\subset\mathbb{P}^{2}_{L_{\mathfrak{q}}}$.
-By Lemma 2.3 these twists are $\mathfrak{q}$-adic approximations to the twists $X_{J[\mathfrak{p}]}^{\pm}(7)\subset\mathbb{P}^{2}_{\mathbb{Q}}$.
+By Lemma 9 these twists are $\mathfrak{q}$-adic approximations to the twists $X_{J[\mathfrak{p}]}^{\pm}(7)\subset\mathbb{P}^{2}_{\mathbb{Q}}$.
 After normalising each equation so that the first non-zero coefficient is equal to $1$, the curves obtained therefore have coefficients in $\mathbb{Q}$ (up to a numerical error).
@@
     We minimise and reduce the models for $\mathcal{X}^{\pm}$ using the algorithm of Elsenhans–Stoll [ES24], which is implemented in Magma as MinRedTernaryForm.
-Remark 3.4.
+Remark 14.
 Computing the matrix $A\in\operatorname{GL}_{3}(L_{\mathfrak{q}})$ must be done with some care in order to control coefficient explosion (and to minimise the $\mathfrak{q}$-adic precision we must carry throughout the calculation).
@@
 The elements $k\zeta_{7}^{i}$ span $L/\mathbb{Q}$ where $0\leq i\leq 6$ and $k$ ranges over the $\operatorname{Gal}(L/\mathbb{Q})$-conjugates of $k_{j}$ for each $1\leq j\leq 7$.
-We then apply Algorithm 3.1 noting Remark 3.2.
+We then apply Algorithm 11 noting Remark 12.
 3.3. Computing the polynomials $g_{i}(t)$
@@
 Let $C/\mathbb{Q}$ be a genus $2$ curve given by a Weierstrass equation $C:y^{2}=f(x)$ whose Jacobian $J/\mathbb{Q}$ has RM by $\mathcal{O}_{D}$, and such that $7$ splits in $\mathcal{O}_{D}$.
-[…]
+Our approach follows that of Fisher [Fis16, Theorem 6.3].
+Using the analytic Jacobian machinery in Magma (in particular the functions AnalyticJacobian and EndomorphismRing) we compute complex approximations to a $\mathfrak{p}$-torsion divisor $\mathfrak{D}=(x_{1},y_{1})+(x_{2},y_{2})-(\infty^{+}+\infty^{-})\in J(\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}\mkern-1.5mu}\mkern 1.5mu)$.
 The model for the Kummer surface $\mathcal{K}$ of $J$ given in [CF96, Chapter 3] and maps $\mathfrak{D}$ to the point $(1:x_{1}+x_{2}:x_{1}x_{2}:\beta_{0})\in\mathcal{K}$ where $\beta_{0}\in\mathbb{Q}(x_{1},x_{2},y_{1},y_{2})$ is the rational function in [CF96, (3.1.4)].
@@
 The prime number $7$ is a norm from $\mathbb{Z}[\sqrt{2}]$ and we may write $(7)=(3+\sqrt{2})(3-\sqrt{2})$.
-Let $[\sqrt{2}]\colon J\to J$ denote the multiplication-by-$\sqrt{2}$-map on $J$.
-The morphism $[\sqrt{2}]$ is a Richelot isogeny and using the approach in [Nic18, Section 5.7] (which is implemented in [Nic]) we determine explicit polynomials giving the morphism $[\sqrt{2}]\colon\mathcal{K}\to\mathcal{K}$ induced by the action of $\sqrt{2}$ on $J$.
-Remark 3.5.
+Let $\sqrt{2}]\colon J\to J$ denote the multiplication-by-$\sqrt{2}$-map on $J$.
+The morphism $\sqrt{2}]$ is a Richelot isogeny and using the approach in [Nic18, Section 5.7] (which is implemented in [Nic]) we determine explicit polynomials giving the morphism $\sqrt{2}]\colon\mathcal{K}\to\mathcal{K}$ induced by the action of $\sqrt{2}$ on $J$.
+Remark 15.
 By interpolation it is not difficult to give an explicit morphism $\mathcal{K}\to\mathcal{K}$ realising the $\sqrt{2}$-action on the Jacobian of the generic member of the generic family of genus $2$ curves $\mathcal{C}/\mathbb{Q}(A,P,Q)$ provided by Bending [Ben98, Ben99].
 We record explicit equations for this (generic) morphism in [Fre].
-Formulae for the multiplication-by-$3$-map $[3]\colon\mathcal{K}\to\mathcal{K}$ are given in [CF96, Section 3.5].
+Formulae for the multiplication-by-$3$-map $3]\colon\mathcal{K}\to\mathcal{K}$ are given in [CF96, Section 3.5].
 Note that $\mathcal{K}[3+\sqrt{2}]\cup\mathcal{K}[3-\sqrt{2}]$ is exactly the set $\{P\in\mathcal{K}:3P=\sqrt{2}P\}$.
@@
 The fundamental discriminants $D<100$ for which this occurs are $D=8$, $29$, $37$, $44$, $53$, $57$, $60$, $65$, $85$, $88$, $92$, and $93$.
-[…]
-[…]
+A generic family of genus $2$ curves $C/\mathbb{Q}$ whose Jacobians have RM by $\mathcal{O}_{8}$ are given by Bending [Ben98, Ben99], who also records many examples of small conductor in [Ben98, Appendix A].
+Bending’s family is given by a triple of parameters $A,P,Q\in\mathbb{Q}$.
+It is simple to search for further examples of small conductor (noting from [Ben98, Section 6.3] that it is often useful to specialise at $P\in\{\pm 1,\pm 1/2,\pm 1/3,\pm 1/5\}$).
+Combining these with the examples found in the LMFDB [LMFDB] we obtain a small (non-exhaustive) database of curves with RM by $\mathbb{Z}[\sqrt{2}]$ and whose Jacobians have conductor $\sqrt{N_{J}}\leq 500\,000$ (these may be found in [Fre]).
+Similar generic families are provided for each $D=8$, $29$, $37$, $44$, and $53$ in [CFM24] building on work of Elkies–Kumar [EK14], who compute the moduli of such curves for all fundamental discriminants $D<100$.
+Some examples of curves with RM by $\mathcal{O}_{D}$ and with small conductor are recorded in [EK14].
+We record a (non-exhaustive) list of such curves with $\sqrt{N_{J}}\leq 500\,000$ in [Fre].
+Note that when $D>17$ the moduli space of curves with RM by $\mathcal{O}_{D}$ is not rational, so examples are sparser than when $D=8$.
 We run the algorithm in Section 3.2 for each curve recorded in [Fre].
-Remark 3.6.
+Remark 16.
 It would be interesting to compute the twists $X_{\mathcal{J}[\mathfrak{p}]}^{\pm}(7)$ for the Jacobian $\mathcal{J}$ of the generic curves $\mathcal{C}/\mathbb{Q}(a,b,c)$ with RM by $\mathcal{O}_{D}$ given in [Ben98, Ben99] and [CFM24] for each $D=8$, $29$, $37$, $44$, and $53$ (i.e., those $D$ where $7$ splits in $\mathcal{O}_{D}$ and for which [CFM24] gives a generic model for a curve by $\mathcal{O}_{D}$).
@@
 We assume that $\mathcal{X}$ is a twist of $X(7)$ (note that this is simple to check by computing Dixmier–Ohno invariants [Dix87, Ohn, Els15] in Magma).
-For the purpose of proving Theorem 1.1 it suffices to consider only the case when $\mathcal{X}$ has a rational point (i.e., it suffices to recall [Fis16, Lemma 6.2], see Lemma 4.2 below).
+For the purpose of proving Theorem 1 it suffices to consider only the case when $\mathcal{X}$ has a rational point (i.e., it suffices to recall [Fis16, Lemma 6.2], see Lemma 18 below).
 In Section 4.2 we note how one may prove that a twist is isomorphic to $X_{M}^{\pm}(7)$ more generally.
@@
 Let $K$ be a field of characteristic zero.
-Suppose that $\mathcal{X}$ has a $K$-rational point which corresponds (through the moduli interpretation in Lemma 2.4) to an elliptic curve $E/K$ (defined up to quadratic twist) with $j$-invariant $j(E)\neq 0,1728,\infty$.
+Suppose that $\mathcal{X}$ has a $K$-rational point which corresponds (through the moduli interpretation in Lemma 10) to an elliptic curve $E/K$ (defined up to quadratic twist) with $j$-invariant $j(E)\neq 0,1728,\infty$.
 In this case, the following lemma reduces the problem of showing that $\mathcal{X}$ is isomorphic to $X_{M}^{\pm}(7)$ to the problem of showing that $X_{E}(7)$ is isomorphic to $X_{M}^{\pm}(7)$.
-Lemma 4.1.
-Let $\mathcal{X}/K$ be a twist of $X(7)$ and suppose that there exsists a point $P\in\mathcal{X}(K)$ with $j(P)\neq 0,1728,\infty$.
+Lemma 17.
+Let $\mathcal{X}/K$ be a twist of $X(7)$ and suppose that there exists a point $P\in\mathcal{X}(K)$ with $j(P)\neq 0,1728,\infty$.
 If $E/K$ is an elliptic curve with $j(E)=j(P)$ then $\mathcal{X}$ is isomorphic to $X_{E}(7)$ over $K$.
@@
 We recall the following lemma of Fisher (based on an argument of Serre using Goursat’s lemma [Ser72, Lemme 8]) which allows us to prove such congruences, up to quadratic twists.
-Lemma 4.2.
+Lemma 18.
 Let $K$ be a number field and let $M$ be a $G_{K}$-module which is isomorphic as an abstract group to $(\mathbb{Z}/p\mathbb{Z})^{2}$ for some $p\geq 5$.
 Suppose that $M$ comes equipped with a ($G_{K}$-equivariant) alternating pairing $M\times M\to\mu_{p}$.
-Let $E/K$ be an elliptic curve with surjective mod $p$ Galois representation, let $x_{M}\colon M\to M/\{\pm 1\}$, and let $x_{E}\colon E\to\mathbb{P}^{1}$ be the quotient by {±$\{\pm 1\}$.
+Let $E/K$ be an elliptic curve with surjective mod $p$ Galois representation, let $x_{M}\colon M\to M/\{\pm 1\}$, and let $x_{E}\colon E\to\mathbb{P}^{1}$ be the quotient by $\{\pm 1\}$.
 If there exist non-identity elements $P\in M$ and $Q\in E[p]$ such that $K(x_{M}(P))=K(x_{E}(Q))$ then there exists a quadratic twist $E^{d}$ of $E$ such that $M\cong E^{d}[p]$.
@@
 One expects that if $X$ and $Y$ are isomorphic (and have no non-trivial automorphisms defined over $\mathbb{Q}$), then for a generic such field the conditions of the lemma hold.
-Lemma 4.3.
+Lemma 19.
 Let $X/K$ and $Y/K$ be (geometrically integral) curves defined over a number field $K$.
@@
 ∎
-Proposition 4.4.
-Conider any of the data in [Fre, data/twists.m] which consists of
+Proposition 20.
+Consider any of the data in [Fre, data/twists.m] which consists of
   1. ()
@@
 Proof.
-If $\mathcal{X}$ has a $\mathbb{Q}$-rational point of small height corresponding to an elliptic curve $E/\mathbb{Q}$, we apply Lemma 4.2.
+If $\mathcal{X}$ has a $\mathbb{Q}$-rational point of small height corresponding to an elliptic curve $E/\mathbb{Q}$, we apply Lemma 18.
 In each case, applying [Ser72, Proposition 19] at several good primes (or using Zywina’s algorithm [Zyw22]) suffices to show that the mod $7$ Galois representation attached to $E/\mathbb{Q}$ is surjective.
@@
 These points correspond to elliptic curves $E_{1}/L_{1}$ and $E_{2}/L_{2}$ whose mod $7$ Galois representations may be seen to be surjective by applying [Ser72, Proposition 19] at several places of good reduction.
-Applying Lemma 4.2 as above shows that $E_{1}$ and $E_{2}$ are $(7,\mathfrak{p})$-congruent to $J$, up to a quadratic twist.
+Applying Lemma 18 as above shows that $E_{1}$ and $E_{2}$ are $(7,\mathfrak{p})$-congruent to $J$, up to a quadratic twist.
 Since the mod $7$ Galois representations of $E_{1}/L_{1}$ and $E_{2}/L_{2}$ are surjective we have $\mathbb{Q}(\mathcal{K}[\mathfrak{p}])\cap L_{1}L_{2}=\mathbb{Q}$, where $\mathcal{K}=J/\{\pm 1\}$ is the Kummer surface of $J$.
@@
 Since the mod $7$ Galois representation attached to $J[\mathfrak{p}]$ is surjective ($J[\mathfrak{p}]$ is isomorphic over $L_{i}$ to a quadratic twist of $E_{i}[7]$ for each $i=1,2$) the element $\tau$ is contained in the centre of $\operatorname{Aut}_{r}(J[\mathfrak{p}])/\{\pm 1\}$ which is isomorphic to $\operatorname{PSL}_{2}(\mathbb{Z}/7\mathbb{Z})$ as an abstract group.
-Therefore $\tau$ is the identity and the claim follows from Lemma 4.3.
+Therefore $\tau$ is the identity and the claim follows from Lemma 19.
 ∎
-5. Proving $(7,\mathfrak{p})$-congruences and Theorem 1.1
-We now prove Theorem 1.1.
-In order to apply visibility we must first show that the pairs $(C,E)$ in Table 1 are in fact $(7,\mathfrak{p})$-congruent (not simply up to quadratic twist, as we proved in Section 4).
-Lemma 5.1.
+5. Proving $(7,\mathfrak{p})$-congruences and Theorem 1
+We now prove Theorem 1.
+In order to apply visibility we must first show that the pairs $(C,E)$ in Table 1.1 are in fact $(7,\mathfrak{p})$-congruent (not simply up to quadratic twist, as we proved in Section 4).
+Lemma 21.
 Let $E/\mathbb{Q}$ be an elliptic curve and let $J/\mathbb{Q}$ be a genus $2$ Jacobian with RM by $\mathcal{O}_{D}$.
 Suppose that $(p)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathcal{O}_{D}$ and that there exists a squarefree integer $d\in\mathbb{Z}$ such that $E^{d}$ and $J$ are $(p,\mathfrak{p})$-congruent.
-Then $d$ is supported on the set of primes consisting of $p$, the bad primes of $E$, and the bad prime of $J$.
+Then $d$ is supported on the set of primes consisting of $p$, the bad primes of $E$, and the bad primes of $J$.
 Proof.
@@
 ∎
-Proposition 5.2.
-For each pair $(E,C)$ of elliptic curve $E/\mathbb{Q}$ and genus 2 curve $C/\mathbb{Q}$ in Table 1 we have a $(7,\mathfrak{p})$-congruence between $E$ and $J=\operatorname{Jac}(C)$ for some choice of $\mathfrak{p}|7$ in $\mathcal{O}_{D}$.
+Proposition 22.
+For each pair $(E,C)$ of elliptic curve $E/\mathbb{Q}$ and genus 2 curve $C/\mathbb{Q}$ in Table 1.1 we have a $(7,\mathfrak{p})$-congruence between $E$ and $J=\operatorname{Jac}(C)$ for some choice of $\mathfrak{p}|7$ in $\mathcal{O}_{D}$.
 Proof.
 Let $\ell\neq 7$ be a good prime for $C$ and $E$.
-By [Fis16, (5.2)] (which follows from [FLS+01, Section 2.1] or [MS93, Lemma 3]) a $(7,\mathfrak{p})$-congruence between $E^{d}/\mathbb{Q}$ and $J/\mathbb{Q}$ gives a congruence modulo $7$
+By [Fis16, (5.2)] (which follows from [FLS+01, Section 2.1] or [MS93, Lemma 3]) a $(7,\mathfrak{p})$-congruence between $E^{d}/\mathbb{Q}$ and $J/\mathbb{Q}$ gives a congruence modulo 77
 (2) $a_{\ell}(E^{d})^{2}-t_{\ell}a_{\ell}(E^{d})+n_{\ell}\equiv 0\pmod{7}$
 where $t_{\ell}=\ell+1-N_{1}$ and $n_{\ell}=(N_{1}^{2}+N_{2})/2-(\ell+1)N_{1}$-$\ell$ where $N_{1}=\#C(\mathbb{F}_{\ell})$ and $N_{2}=\#C(\mathbb{F}_{\ell^{2}})$.
-Testing (2) on the divisors $d$ of the product $7$ and the bad primes of $E$ and $C$ shows that $E^{d}$ and $J$ are not $(7,\mathfrak{p})$-congruent for any $d\neq 1$ (by Lemma 5.1).
-By 4.4 $E$ and $J$ are $(7,\mathfrak{p})$-congruent up to quadratic twist since $E$ corresponds to a point on one of the twists $X_{J[\mathfrak{p}]}^{\pm}(7)$.
+Testing (2) on the divisors $d$ of the product of $7$ and the bad primes of $E$ and $C$ shows that $E^{d}$ and $J$ are not $(7,\mathfrak{p})$-congruent for any $d\neq 1$ (by Lemma 21).
+By 20 $E$ and $J$ are $(7,\mathfrak{p})$-congruent up to quadratic twist since $E$ corresponds to a point on one of the twists $X_{J[\mathfrak{p}]}^{\pm}(7)$.
 It therefore follows that $E$ and $J$ are $(7,\mathfrak{p})$-congruent for some choice of $\mathfrak{p}$ dividing $7$.
 ∎
-Using the congruences supplied by 5.2 we now prove Theorems 1.1 and 1.5 by applying [Fis16, Theorem 2.2].
-Proof of Theorems 1.1 and 1.5.
+Using the congruences supplied by 22 we now prove Theorems 1 and 5 by applying [Fis16, Theorem 2.2].
+Proof of Theorems 1 and 5.
 This follows from [Fis16, Theorem 2.2], as we detail below.
-Let $E/\mathbb{Q}$ and $J/\mathbb{Q}$ be one of the pairs of elliptic curve and genus $2$ Jacobian from Theorem 1.1 or 1.5.
+Let $E/\mathbb{Q}$ and $J/\mathbb{Q}$ be one of the pairs of elliptic curve and genus $2$ Jacobian from Theorem 1 or 5.
 We check that in each case $J$ is geometrically simple by applying the condition in [CF96, Section 14.4] and [Sto95].
 The $7$-torsion subgroups of $E(\mathbb{Q})$ and $J(\mathbb{Q})$ are trivial.
-The rank of $E/\mathbb{Q}$ is $2$ and the rank of $J/\mathbb{Q}$ is 00 (the rank of $J/\mathbb{Q}$ is bounded using $2$-descent, which is implemented as RankBounds in Magma).
-For each discriminant $D$ appearing in Theorems 1.1 and 1.5 the prime $7$ not only splits in $\mathcal{O}_{D}$, but $7=\operatorname{Nm}\eta$ for some $\eta\in\mathcal{O}_{D}$.
+The rank of $E/\mathbb{Q}$ is $2$ and the rank of $J/\mathbb{Q}$ is 0 (the rank of $J/\mathbb{Q}$ is bounded using $2$-descent, which is implemented as RankBounds in Magma).
+For each discriminant $D$ appearing in Theorems 1 and 5 the prime $7$ not only splits in $\mathcal{O}_{D}$, but $7=\operatorname{Nm}\eta$ for some $\eta\in\mathcal{O}_{D}$.
 In particular the isogeny $\mathfrak{p}$ is equal to the multiplication-by-$\eta$-map on $J$, and $J(\mathbb{Q})/\mathfrak{p}J(\mathbb{Q})=0$.
-By 5.2 the elliptic curve $E$ is $(7,\mathfrak{p})$-congruent to $J$.
+By 22 the elliptic curve $E$ is $(7,\mathfrak{p})$-congruent to $J$.
 The abelian varieties $E$ and $J$ have good reduction at $7$, so by [Fis16, Theorem 2.2] it suffices to show that the Tamagawa numbers of $E/\mathbb{Q}$ and $J/\mathbb{Q}$ are coprime to $7$.
 We compute the Tamagawa numbers of $E/\mathbb{Q}$ using Magma.
-Except for the Jacobian of conductor $3200^{2}$ in Table 1, for each bad prime $p$ of $J$ one may check that the order of the geometric component group of $J$ at $p$ is coprime to $7$ using Liu’s genus2reduction in SageMath and Donnelly’s Magma functions RegularModel and ComponentGroup.
-For the Jacobian of conductor $3200^{2}$ in Table 1 the computation of the Tamagawa number of $J/\mathbb{Q}$ at $2$ was carried out in the appendix to [KS23] (where it is shown that the Tamagawa number is $1$).
+Except for the Jacobian of conductor $3200^{2}$ in Table 1.1, for each bad prime $p$ of $J$ one may check that the order of the geometric component group of $J$ at $p$ is coprime to $7$ using Liu’s genus2reduction in SageMath and Donnelly’s Magma functions RegularModel and ComponentGroup.
+For the Jacobian of conductor $3200^{2}$ in Table 1.1 the computation of the Tamagawa number of $J/\mathbb{Q}$ at 22 was carried out in the appendix to [KS25] (where it is shown that the Tamagawa number is 11).
 ∎
-6. Evidence towards 1.6
+6. Evidence towards 6
 Consider the genus $2$ curve $C/\mathbb{Q}$ with LMFDB label 385641.a.385641.1 and Weierstrass equation
@@
 The Jacobian $J/\mathbb{Q}$ of $C$ has RM by $\mathbb{Z}[\sqrt{2}]$.
-In [KS23, A.3] it is noted that the Birch and Swinnerton-Dyer conjecture predicts $|\Sha(J^{-11}/\mathbb{Q})|=7^{2}$.
-By 4.4, for some choice of factorisation $(7)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathbb{Z}[\sqrt{2}]$ we have models
+In [KS25, A.3] it is noted that the Birch and Swinnerton-Dyer conjecture predicts $|\Sha(J^{-11}/\mathbb{Q})|=7^{2}$.
+By 20, for some choice of factorisation $(7)=\mathfrak{p}\bar{\mathfrak{p}}$ in $\mathbb{Z}[\sqrt{2}]$ we have models
 $X_{J[\mathfrak{p}]}^{\pm}(7):$ $\;-2x_{0}^{4}+39x_{0}^{3}x_{1}+11x_{0}^{3}x_{2}-42x_{0}^{2}x_{1}^{2}-18x_{0}^{2}x_{1}x_{2}+20x_{0}x_{1}^{3}-6x_{0}x_{1}^{2}x_{2}$
@@
 We were unable to find rational points on any of these curves, except on $X_{J[\mathfrak{p}]}^{\mp}(7)$ where we find exactly one point which corresponds to the elliptic curve $E/\mathbb{Q}$ with LMFDB label 1242.m1 and Weierstrass equation $y^{2}+xy+y=x^{3}-x^{2}-1666739x-2448131309$.
-Using the argument in 5.2 it can be shown that $E$ and $J$ are $(7,\mathfrak{p})$-congruent.
-However, the quadratic twist of $E$ by $-11$ has trivial Mordell-Weil group, so cannot be used to visualise the (conjectural) non-trivial elements of $\Sha(J^{-11}/\mathbb{Q})[7]$.
+Using the argument in 22 it can be shown that $E$ and $J$ are $(7,\mathfrak{p})$-congruent.
+However, the quadratic twist of $E$ by 11-11 has trivial Mordell-Weil group, so cannot be used to visualise the (conjectural) non-trivial elements of $\Sha(J^{-11}/\mathbb{Q})[7]$.